### v1

# A Comprehensive Survey of Deep Transfer Learning for Anomaly Detection in Industrial Time Series: Methods, Applications, and Directions

Abstract
Automating the monitoring of industrial processes has the potential to enhance efficiency and optimize quality by promptly detecting abnormal events and thus facilitating timely interventions. Deep learning, with its capacity to discern non-trivial patterns within large datasets, plays a pivotal role in this process. Standard deep learning methods are suitable to solve a specific task given a specific type of data. During training, the algorithms demand large volumes of labeled training data. However, due to the dynamic nature of processes and the environment, it is impractical to acquire the needed data for standard deep learning training for every slightly different case anew. Deep transfer learning offers a solution to this problem. By leveraging knowledge from related tasks and accounting for variations in data distributions, this learning framework solves new tasks even with little or no additional labeled data. The approach bypasses the need to retrain a model from scratch for every new setup and dramatically reduces the labeled data requirement. This survey provides an in-depth review of deep transfer learning, examining the problem settings of transfer learning and classifying the prevailing deep transfer learning methods. Moreover, we delve into applying deep transfer learning in the context of a broad spectrum of time series anomaly detection tasks prevalent in primary industrial domains, e.g., manufacturing process monitoring, predictive maintenance, energy management, and infrastructure facility monitoring. We conclude this survey by underlining the challenges and limitations of deep transfer learning in industrial contexts. We also provide practical directions for solution design and implementation for these tasks, leading to specific, actionable suggestions.
Keywords deep transfer learning  $\cdot$ time series analysis  $\cdot$ anomaly detection  $\cdot$ manufacturing process monitoring  $\cdot$ predictive maintenance
1 Introduction
The fourth industrial revolution — Industry 4.0 (Kagermann et al., 2011), that is characterized by increasing efficiency through the digitization of production, automation, and horizontal integration across companies (Roblek et al., 2016), and the advent of connected cyber-physical systems – referred to as internet of things (Wang et al. 2015; Jeschke et al. 2017; Dalenogare et al. 2018), increases the need for autonomous and intelligent process monitoring. This can be exemplified by the use case of a smart factory in which industrial processes are transformed to be more flexible, intelligent, and dynamic (Kagermann, 2017), or the use case of decentralized energy production with wind and solar (Abir et al., 2021). In these examples, AI-powered anomaly detection integrates the analysis of time series data to detect unusual patterns in the recorded data. By identifying parameters that fall outside a window of normal operation, operators can trigger interventions and adjustments to ensure high product quality and safe operations. To achieve this, physical properties such as pressure or temperature are monitored and analyzed in real-time. Changes in these variables capture drifting and abrupt faults caused by process failures or malfunctions (Park et al., 2020). The production process must adapt quickly to changes in production and the environment to meet the requirements for flexibility and dynamics. Further use cases exist in such diverse areas as manufacturing monitoring including automatic quality control, predictive maintenance of goods and services, infrastructure monitoring of e.g. building energy systems or power plants, digital agriculture, petrochemical process optimization, computer network intrusion detection, or aircraft flight monitoring, to name a few.
Artificial Intelligence, in particular deep learning, provides competent frameworks to automate intelligent monitoring in order to provide valuable assistance to operators and high-level control systems. Leveraging the power of deep learning, formative features of the data – technically referred to as representations (Bengio et al., 2013) – can be captured in a machine-learned model and thereby enable a detailed understanding of variations in standard operations.
However, in non-trivial and non-stationary conditions, the task or the underlying data may change. For example, the monitoring system of a milling machine may in one instance be tasked to detect a blunt tool based on the vibration and in another instance – using the same vibration measurements – to detect insufficient cooling lubricant. Knowledge acquired to solve one task in one setting with a given tool, machined part, and type of machine may be transferred to solve the same or similar task in a setting with a different tool, machined part, or type of machine. Slowly changing conditions (drifts), abrupt mode changes (for instance due to tool change), and new tasks (such as the detection of another failure mode) may require adjustments to the machine-learned model. In these cases, it is desirable to adjust the analysis model without retraining from scratch, as it is costly or impractical to acquire sufficient training data to learn the full manifold (Maschler et al., 2022).
Transfer learning is a machine learning framework to achieve this (Maschler and Weyrich, 2021; Zhuang et al., 2021; Pan and Yang, 2010; Tan et al., 2018; Yosinski et al., 2014). As depicted in Figure 1, data and algorithms from a related task may be leveraged in a new one. By accounting for changes in data distributions and tasks, or leveraging existing models, knowledge learned from related tasks can be used to improve performance on new tasks instead of retraining a model for each individual application from scratch. This transfer-learning-boosted modeling forms the basis for identifying anomalies that deviate from established patterns in a non-trivial manner without full re-training.
Figure 1: Transfer learning is useful when changes in production take place and sufficient data for full retraining is not available as shown here for a hypothetical production of two types of gears. In the production of gear A, a lot of data is available to train a deep learning model that helps improve production. In the production of gear B, data availability is more limited, and the traditionally trained deep learning model fails to improve production. With suitable methods, however, data and algorithms acquired during production of gear A can be leveraged to support improving the production of gear B, because the data and tasks in the production of both gears are related.
Deep transfer learning (Tan et al., 2018; Yu et al., 2022) extends the transfer learning paradigm by leveraging deep learning models. In industrial contexts, it ensures optimal production even as production conditions shift. This dynamic adaptability is key in maintaining the effectiveness of anomaly detection systems in the dynamic environment that characterizes industrial applications including the broad categories of manufacturing process monitoring, predictive maintenance, energy management, and infrastructure facility monitoring as detailed in Section 4.
In this survey, we review the foundations of deep transfer learning to equip the reader with working knowledge of the main principles and intuition for ideas. Further, we provide a comprehensive overview of the current state of the art of deep transfer learning approaches for time series anomaly detection for industrial applications. Our main contribution is a systematic review of research work on real-world industrial applications. For these, we discuss potential, challenges, and limitations and give directions for future work and potential. The paper is organized as follows: First, we introduce a taxonomy of transfer learning problem settings and further categorization of deep transfer learning approaches (Section 2). Then, we describe the task of anomaly detection in time series (Section 3) in selected industrial applications (Section 4). To conclude, we discuss current challenges, limitations, and future research directions (Sections 56) in the field.
2 Deep transfer learning
Transfer learning in a deep learning setting aims to increase the efficiency, performance, and generalization of deep learning models by transferring knowledge from one data set and task to a new one. This eliminates the need to train a deep learning model from scratch, which in turn reduces the amount of necessary data and compute required to solve a new task or new data domain. In either case, knowledge is transferred from a source to a target domain, as defined below. The transfer learning problem settings can be categorized as inductive or transductive transfer depending on the data and task conditions, while we categorize deep learning-based transfer learning approaches into instance transfer, parameter transfer, mapping transfer and domain-adversarial transfer. We illustrate them by using two intuitive examples in Figure 2, with more details being elaborated in the following sections.
Figure 2: Taxonomy of transfer learning problem settings (left; see Section 2.1 for the definition of terms) and corresponding examples using deep transfer learning approaches (right). On the left, we classify transfer learning problems as inductive or transductive transfer settings, mainly depending on the label availability and task similarities. Correspondingly, we provide two examples using deep transfer learning methods: In the inductive transfer setting, we collect time series data from screw production and wrench production. Labeled screw data (A1) is used to detect collective anomalies (a set of data points behaving differently compared to the entire time series (Choi et al., 2021; Chalapathy and Chawla, 2019), further explained in Section 3). Then, parameter transfer (Section 2.2.2) is applied to transfer knowledge by fine-tuning the pre-trained model from labeled screw data to detect point anomalies (further explained in Section 3) on labeled wrench data (A2). For the transductive transfer setting in the lower panel, we present a different situation for contextual anomaly detection (further explained in Section 3). In this case, we have two datasets, B1 and B2, analyzed using the same model. However, the data in B2 significantly differs in appearance from the data in B1. To address this issue, instance transfer (further explained in Section 2.2.1) is used. Through this learning process, the data in B2 is transformed in a way that makes it compatible with the model which has been trained exclusively on data from B1. Transfer learning, in this case, is thus achieved by adapting the data to fit the model through domain adaptation rather than adjusting the model to fit new data.
2.1 Transfer learning problem definition
Domain $\mathcal{D}$ includes the domain feature space $\mathcal{X}$ and marginal data distribution $P(X)$ as $\mathcal{D}=\{\mathcal{X},P(X)\}$, where $X$ is the domain data, $X=\left\{x_{1},\ldots,x_{n}\right\}\in\mathcal{X}$. Similarly, a learning task is defined as $\mathcal{T}=\{\mathcal{Y},f_{\mathcal{T}}(\cdot)\}$, where $\mathcal{Y}$ denotes the task space and usually represents class label. For anomaly detection tasks, $\mathcal{Y}$ is the set of the two classes “normal” and “abnormal”. The function $f_{\mathcal{T}}(\cdot)$ can be used to predict the corresponding label of a new instance $x_{i}$. The objective predictive function $f_{\mathcal{T}}(\cdot)$ can be learned from domain data and can be interpreted as a form of conditional probability. Thus, the learning task can be rewritten as $\mathcal{T}=\{\mathcal{Y},P(Y|X)\}$, where $P(Y|X)$ is used as a likelihood measure to determine how well a given data set $X$ fits with a corresponding class label set $Y$.
We largely follow the definition of transfer learning by Pan and Yang (2010) and Zhuang et al. (2021). Given a source domain $\mathcal{D}_{S}$ and learning task $\mathcal{T}_{S}$, as well as a target domain $\mathcal{D}_{T}$ and learning task $\mathcal{T}_{T}$, transfer learning aims to improve the performance of the predictive function $f_{\mathcal{T}}(\cdot)$ in $\mathcal{D}_{T}$ by transferring knowledge from $\mathcal{D}_{S}$ and $\mathcal{T}_{S}$, where $\mathcal{D}_{S}\neq\mathcal{D}_{T}$ and/or $\mathcal{T}_{S}\neq\mathcal{T}_{T}$. Usually, the size of source dataset is much smaller than target dataset.
This definition of transfer learning can be broadened, i.e., the target task can profit from multiple source domains. Transfer learning is thus the idea of making the best use of related source domains to solve new tasks. In contrast, traditional machine learning (ML) methods learn each task separately from scratch, and each respective model can only be applied to the corresponding task.
We define a taxonomy of transfer learning problems settings as shown in Figure 2 mainly depending on the label availability in the two domains to be easily applicable to the requirements of a case at hand (compare different definitions for other purposes in the literature Pan and Yang (2010); Tan et al. (2018); Zhuang et al. (2021)).
We differentiate it into inductive and transductive transfer learning (Pan and Yang, 2010) [^1] In this survey, we do not consider unsupervised learning scenarios since either source labels or target labels are provided for most industrial applications.. Inductive transfer learning is applied when the target task is different from the source task, i.e., $\mathcal{T}_{S}\neq\mathcal{T}_{T}$ (meaning that $\{\mathcal{Y_{S}}\neq\mathcal{Y_{T}}\}$ or $\{P(Y_{S}|X_{S})\neq P(Y_{T}|X_{T})\}$). The conditional probability distribution is induced with labeled training data in the target domain (Torrey and Shavlik, 2009). A corresponding example is illustrated as Scenario A in Figure 2, where the learning tasks are different and the goal of transfer learning is to recognize point anomaly from the collective anomaly task. Related areas of inductive transfer learning are multi-task learning (Caruana, 1997; Ruder, 2017) and sequential learning, depending on whether tasks are learned simultaneously or sequentially.
Transductive transfer learning is applied when the source and target tasks are the same, while the source and target domain are different, i.e., $\mathcal{T}_{S}=\mathcal{T}_{T}$ and $\mathcal{D}_{S}\neq\mathcal{D}_{T}$ (meaning that $\{\mathcal{X_{S}}\neq\mathcal{X_{T}}\}$ or $\{P(X_{S})\neq P(X_{T})\}$. A subcategory is domain adaptation (Kouw and Loog, 2019) when the feature space of source and target data are the same but the corresponding marginal distributions are different (i.e., $\{\mathcal{X_{S}=X_{T}}\}$ and $\{P(X_{S})\neq P(X_{T})\}$). Scenario B in Figure 2 is an example of transductive transfer learning where the learning tasks are identical, and the goal of transfer learning is to recognize contextual anomalies in an unlabelled data set.
Other learning paradigms closely related to transfer learning are listed below:
Multi-task learning is a machine learning technique where a single model is trained on multiple tasks simultaneously. The idea is to improve the performance of the model by learning a shared representation that captures the features between all tasks.
Continuous learning (Parisi et al., 2019) is a learning process where the model continuously learns new tasks from previous tasks over time without forgetting how to solve previous tasks. To some extent, continuous learning can be seen as a sequential transfer learning process, with the constraint to preserve the performance on the previous tasks which leads to an accumulation of knowledge over time.
Few-shot learning (Wang et al., 2020) is a type of machine learning where a model can learn and perform well on a new task with only a limited number of labeled samples. In extreme cases, the model can learn with one label (Fei-Fei et al., 2006) and without any label (Lampert et al., 2009). While transfer learning usually involves reusing the model from relevant tasks and continuing training on the target dataset.
Meta-learning (Finn et al., 2017; Hospedales et al., 2022) is a machine learning technique that focuses on the learning process. It is known as “learning to learn”. For meta-learning, models are trained on a different set of tasks instead of a set of data in the traditional machine learning setting. In this sense, meta-learning can be seen as a form of transfer learning because it involves transferring knowledge from task to task.
Knowledge distillation (Gou et al., 2021) effectively learns a small model trained to mimic the behavior of a larger, more complex model. The knowledge learned by the larger model can be transferred to the smaller model, which can then be used for the target task, e.g., on a less powerful edge device.
Self-supervised learning (Liu et al., 2023; Bai et al., 2021) involves training a model to predict some aspect of the input data without any external supervision. The learned representations can be used for various downstream tasks, including those that involve transferring knowledge from one domain to another.
2.2 Deep transfer learning approaches
Since deep neural networks (DNNs) can learn useful feature representations from large amounts of data through back-propagation (Bengio et al., 2013), they have been widely adopted for tackling complex problems, which involve large-scale and high-dimensional data, also in practice (Schmidhuber, 2015; Lukic et al., 2016; Stadelmann et al., 2018; Schmidhuber, 2022). Deep transfer learning methods implement transfer learning principles within DNN and, among other things, enable deep learning based analysis pipelines to be applied to new datasets. On a high level, deep transfer learning approaches can be divided into data-driven and model-driven ones. Data-driven approaches focus on transferring knowledge by transforming and adjusting the data instances. Model-driven approaches leverage DNNs to develop domain-invariant features by reducing the feature discrepancy between source and target domain data and then transferring generalized knowledge to new tasks. Following the taxonomy of (Tan et al., 2018; Pan and Yang, 2010; Shi et al., 2022), we divide deep transfer learning approaches further into $4$ categories: instance transfer, parameter transfer, mapping transfer, and domain-adversarial transfer, as illustrated in Table 1. Instance transfer and mapping transfer are data-driven approaches, parameter transfer is a model-driven approach, while domain-adversarial transfer is a combination of both.
Table 1: Overview of deep transfer learning approaches with references.
Deep transfer learning approach Short description References
Instance transfer
Augmenting target data by transforming data instance from the source domain
He et al., 2022; Amirian et al., 2021; Wang et al., 2018
Parameter transfer
Transfer learned parameters of a pre-trained model from source domain and adapting the model for target domain
Bommasani et al.,2022; Devlin et al.,2019; Vaswani et al.,2017; Dou et al.,2022; Alexandr et al.,2021; Yosinski et al.,2014; Guo et al.,2019; Sager et al.,2022; Brown et al.,2020
Mapping transfer
Reducing feature discrepancies between source and target domains by minimizing the distance between mapped features in the latent space
Zhuang et al., 2021; Tzeng et al., 2014; Long et al., 2017, 2015; Zhang et al., 2015; Venkateswara et al., 2017
Domain-adversarial transfer
Extracting an indiscriminative feature representation between source and target domain through adversarial training
Tzeng et al., 2015; Ganin et al., 2017; Ajakan et al., 2015; Tzeng et al., 2017
2.2.1 Instance transfer
The intuition of instance transfer is that although source and target domains differ, it is still possible to transform and reuse source data together with a few labeled target samples. A typical approach is to re-create some labeled data from the source domain. He et al. (2022) propose an instance-based deep transfer learning model with attention mechanism to predict stock movement. They first create new samples from the source dataset that are similar to the target samples by using attention weights, and then train on the created samples and target training samples for prediction tasks. Amirian et al. (2021) introduce an innovative instance transfer method for domain adaptation. They propose an effective auto-encoder model with a pseudo-label classifier to reconstruct new data instances that obtain general features across different datasets for medical image analysis. Taking another avenue, Wang et al. (2018) exclude the source data that have a bad impact on training target data. Specifically, they choose a pre-trained model from a source domain, estimate the impact of all training samples in the target domain, and remove samples that lower the performance of the model.
2.2.2 Parameter transfer
Parameter transfer adapts learned parameters of a pre-trained model to a new model. This assumes that DNNs can get similar feature representations from similar domains. Thus, through transferring parts of the DNN layers together with pre-trained parameters and/or hyperparameters, the pre-trained model is used as a base model to further train on target domain data and solve different learning tasks. Particularly, parameter transfer has gained popularity in computer vision and natural language processing, where large models are pre-trained on large datasets (Bommasani et al., 2022). In natural language processing, BERT (Devlin et al., 2019) and GPT-3 (Brown et al., 2020) that are based on the transformer architecture (Vaswani et al., 2017) can be fine-tuned for a variety of natural language processing tasks, including content generation (Dou et al., 2022), language translation (Sun et al., 2021), question answering (Glass et al., 2020), and summarization (Alexandr et al., 2021). Yosinski et al. (2014) investigated the general transferability of DNNs. Experiment results show that transferring features from source to target domain leads to improved generalization in networks compared to those trained solely on the target dataset. Unlike the typical way of fine-tuning a pre-trained model, Guo et al. (2019) propose the adaptive fine-tuning approach SpotTune to find the optimal fine-tuning strategy for the target task. Specifically, a policy network is used to make routing decisions on whether to pass the target instance through the pre-trained model. The results show SpotTune is effective in most cases by using a hybrid of parameter and instance transfer. Sager et al. (2022) propose an unsupervised domain adaptation for vertebrae detection in 3D CT volumes by transferring knowledge across domains during each batch of the training process.
2.2.3 Mapping transfer
Mapping transfer refers to learning a related feature representation for the target domain by feature transformation, which includes feature alignment, feature mapping, and feature encoding (Zhuang et al., 2021). The goal is to reduce feature discrepancies between source and target domains by minimizing the distance between the distribution of mapped features in the latent space. There are various criteria to measure the distribution difference, including Wasserstein distance (Shen et al., 2018), Kullback-Leibler Divergence (Dai et al., 2007), etc. Among them, Maximum Mean Discrepancy (MMD) (Tzeng et al., 2014) is most frequently adopted in mapping transfer from the surveyed papers. The MMD is calculated as the difference between the mean embeddings of the samples in a reproducing kernel Hilbert space associated with a chosen kernel function. Added to the target loss function, it measures the difference between two probability distributions and serves as a powerful tool for comparing the similarity of complex, high-dimensional datasets using a wide variety of kernel functions.
Some previous work focusing on transferred feature extraction/dimensionality reduction using MMD has been done. Long et al. (2017) base their Joint Adaptation Network on MMD, in which the joint distributions of multiple domain-specific layers across domains are aligned. In addition, an adversarial training version was adopted to make distributions of the source and target domains more distinguishable. Similarly, Long et al. (2015) adopted multi-layer adaptation and proposed Deep Adaptation Networks (DAN). In DAN models, the first three convolutional layers are used to extract general features. For the last three layers, multi-kernel MMD is used to bridge the cross-domain discrepancy and learn transferable features. Zhang et al. (2015) also based on MMD and proposed a Deep Transfer Network in which two types of layers are used to obtain domain invariant features across domains. The shared feature extraction layers learn a shared feature subspace between the source and the target samples, and the discrimination layer is then used to match conditional distributions by classifier transduction. Venkateswara et al. (2017) proposed Deep Adaptation Hash network, which is fine-tuned from the VGG-F (Chatfield et al., 2014) network. Multi-kernel MMD loss trains the Deep Adaptation Hash to learn feature representations that align the source and target domains.
2.2.4 Domain-adversarial transfer
Inspired by Generative Adversarial Networks (GANs) (Goodfellow et al., 2014; Schmidhuber, 2020), the goal here is to extract a transferable feature representation that is indiscriminative between source and target domain through adversarial training. Adversarial transfer mainly focuses on domain adaptation problems. Tzeng et al. (2015) adopt a domain confusion loss across the source and target domains to learn a domain invariant representation. Ganin et al. (2017) propose a new domain adaptation architecture by adding a domain classifier after feature extraction layers. A gradient reversal layer is used to ensure the similarity of the feature distributions over source and target domains. Ajakan et al. (2015) propose a domain adversarial DNN in which a domain regressor is applied to learn a domain invariant feature representation. Tzeng et al. (2017) use an unsupervised domain adaptation method that combines adversarial learning with discriminative feature learning.
3 Time series analysis for industrial processes
Time series analysis encompasses statistical techniques to analyze and interpret sequential temporal data. In the context of industrial processes, time series analysis plays a crucial role in automating monitoring and controlling the efficiency, quality, and performance of these processes. Specifically, the analysis of time series data can be used for anomaly detection, forecasting, process control, performance assessment, and maintenance scheduling to increase the efficiency of the process.
Anomaly detection
According to Hawkins (1980), an outlier is as an observation that deviates significantly from other observations in a way that it is likely that it was generated by a different mechanism. In this survey, we focus on time series data collected from machine sensor readings in the context of industrial applications, either univariate (only one variable is recorded over time) or multivariate (several simultaneously recorded measurements). Time series anomalies might occur for various reasons, including internal factors (e.g., temporary sensor error, machinery malfunction) and external factors (e.g. human error, ambient temperature). They can be divided into three categories (Choi et al., 2021; Chalapathy and Chawla, 2019): point anomalies, contextual anomalies, and collective anomalies. Point anomalies are isolated samples that deviate significantly from the normal behavior of that time series, which can be seen on the left of Fig. 3, e.g., a sudden spike in a pressure reading from a manufacturing machine sensor. These point anomalies can be caused by temporal sensor error, human error, or abnormal machinery operations. Contextual anomalies represent data points that deviate from normal ones only in their current context, and an example can be seen in the middle of Fig. 3. Collective anomalies are a set of data points that in their entirety (but not individually) are abnormal with respect to the entire time series, as shown on the right of Fig. 3.
Figure 3: Three time series anomaly types. Gray lines represent recorded time series signals, and dashed green lines are a priori set thresholds of normal operations. The red dots and the red line indicate anomalies. Point anomalies are single values that fall outside of a pre-set range (left panel). Contextual anomalies are samples that deviate from current context (middle panel). Collective anomalies are defined as a series of data points that all fall within the range of operation but jointly are not expected (right panel).
Challenges regarding detecting time series anomalies persist due to two specific properties: First, the complexity of time series data. As the automation level of industrial processes and the complexity of industrial systems increases, univariate time series data become insufficient and inefficient in representing any industrial process in its entirety. Hence, more sensors are installed to monitor the whole process, making it necessary in turn to detect anomalies from multivariate time series, which poses particular challenges since it requires consideration of temporal dependencies and relationships between variables and modalities. Second, the dynamic variability of industrial processes can pose difficulties in detecting anomalies due to fluctuations in the process caused by varying input or environmental conditions such as material, temperature, pressure, and humidity, which lead to domain shifts.
Process automation
The tasks described above are combined to automate industrial processes. For example, after the detection of an anomaly, another model that captures the relationship between time course and different failure modes or drifts may be exploited for predictive maintenance. For example, in injection molding process monitoring, anomaly detection models are used to analyze recorded sensor data from injection molding machines to detect bad parts and identify the root cause of anomalies (Tercan et al., 2018). There are two basic ways to detect anomalies: for supervised anomaly detection, labels (normal/abnormal) are needed per time series to build a binary classifier (Görnitz et al., 2013). For unsupervised anomaly detection, an anomaly score or confidence value that is conditioned purely on normal data can be used to differentiate abnormal from normal instances (Zhang et al., 2023; Audibert et al., 2020; Stadelmann et al., 2019).
4 Industrial applications
4.1 Overview
Currently, deep transfer learning approaches are popular in the field of computer vision and natural language processing because of the large available datasets. This popularity is not as pronounced for industrial time series data, likely due to the lack of publicly available data and the domain-specific differences between such data, making the field a less easy target for general gains. Fortunately, in recent years a growing number of deep transfer learning approaches have been applied in industry to solve anomaly detection tasks, such as fault diagnosis (Li et al., 2022), quality management (Ma et al., 2019), manufacturing process monitoring (Tercan et al., 2018), network/software security (Rosenberg et al., 2018), and infrastructure monitoring (Pan et al., 2023). These can be mapped onto the core industrial domains of manufacturing process and infrastructure monitoring, predictive maintenance, and energy management. Table 2 presents a compact comparison of the related works using deep transfer learning approaches to solve these tasks.
Figure 4 illustrates the quantity structure of the connections between industrial applications and the deep transfer learning approaches based on our literature survey. The Sankey diagram shows every path that connects the four dimensions of the methodology-problem-landscape within the surveyed literature. The broader the path is, the more papers are related to that element. The goal is to give an overview of how deep transfer learning is applied to industrial problems in the recent literature and specifically show with these four dimensions: (1) which deep transfer learning approaches are actually used in practice; (2) what the main industrial domains for time series anomaly detection are; (3) what deep transfer learning category these domains belong to; and (4) what labels are available in source and target domain.
Key observations from Figure 4 are: Regarding deep transfer learning approaches, parameter transfer is much more frequently used than any other deep transfer learning approach across all surveyed industrial applications since fine-tuning a pre-trained model on target data is more straightforward to implement by taking advantage of the pre-trained model on the source dataset and usually without fundamental modification on the model architecture. It is noteworthy that instance transfer and adversarial transfer do not appear in the diagram. Apparently, these two deep transfer learning approaches are not considered the optimal choice for respective time series anomaly detection tasks in the industry. The difficulty lies in implementing and training these scarcely researched approaches in the industrial field, as indicated by the findings. Regarding industrial applications, hybrid approaches of parameter and mapping transfer can be seen in predictive maintenance. Regarding deep transfer learning categories, most industrial applications use inductive TL, indicating they focus on leveraging labeled source and target data to solve the target task, i.e., use supervised learning.
Table 2: A compact overview of industrial applications that used deep transfer learning for time series anomaly detection.
Reference
Industrial task
Industrial domain
Deep transfer learning approach
Transfer learning problem setting
Deep learning framework
Source data type
Source label
Target label
Maschler et al.,2021
Industrial metal forming anomaly detection
Predictive maintenance
Parameter transfer
Inductive
CNN
Multiple sources
Wen and Keyes,2019
Monitoring systems Anomaly detection
Predictive maintenance
Parameter transfer
Inductive
U-Net
Multiple sources
Xu et al.,2019
Car body-side production line fault diagnosis
Predictive maintenance
Parameter transfer
Inductive
SAE
Multiple source
Mao et al.,2020
Rotation bearings fault detection
Predictive maintenance
Mapping transfer
Transductive
Auto-encoder
Multiple sources
Wang et al.,2021
Industrial control systems anomaly detection
Predictive maintenance
Parameter transfer
Inductive
ResNet8
Single source
Canizo et al.,2019
Service elevator fault detection
Predictive maintenance
Parameter transfer
Inductive
CNN+RNN
Multiple sources
Yao et al.,2022
Nuclear power plants fault detection
Predictive maintenance
Parameter transfer
Inductive
CNN
Multiple sources
Li et al.,2023
Building energy systems fault diagnosis
Predictive maintenance
Parameter transfer
Inductive
CNN
Multiple source
Serradilla et al.,2021
Press machine production prediction
Predictive maintenance
Parameter transfer
Inductive
CNN
Single source
Zgraggen et al.,2021
Wind turbine fault detection
Predictive maintenance
Parameter transfer
Inductive
CNN
Single source
Zabin et al.,2023
Industrial machine operating fault detection
Predictive maintenance
Parameter transfer
Inductive
CNN+LSTM
Single source
Liao et al.,2021
Machine turning operations classification
Manufacturing process monitoring
Parameter transfer
Inductive
VGG-19, ResNet
Single source
Lockner and Hopmann,2021
Injection molding process quality control
Manufacturing process monitoring
Parameter transfer
Inductive
FCN
Multiple sources
Tercan et al.,2018
Injection molding process quality control
Manufacturing process monitoring
Parameter transfer
Inductive
FCN
Single source
Tercan et al.,2019
Injection molding process anomaly detection
Manufacturing process monitoring
Parameter transfer
Inductive
FCN
Multiple sources
Lockner et al.,2022
Injection molding process anomaly detection
Manufacturing process monitoring
Parameter transfer
Inductive
FCN
Multiple sources
Gellrich et al.,2021
Aluminum gravity die casting quality prediction
Manufacturing process monitoring
Parameter transfer
Inductive
FCN
Single source
Abdallah et al.,2021
Digital agriculture and smart manufacturing systems anomaly detection
Manufacturing process monitoring
Parameter transfer
Inductive
LSTM
Single source
Abdallah et al.,2023
Manufacturing testbeds anomaly detection
Manufacturing process monitoring
Parameter transfer
Inductive
LSTM, RNN
Single source
Maschler et al.,2021
Industrial metal (pump) forming anomaly detection
Manufacturing process monitoring
Parameter transfer
Inductive
LSTM
Single source
Panjapornpon et al.,2023
Petrochemical production process anomaly detection
Energy saving
Parameter transfer
Inductive
LSTM, CNN, FCN
Single source
Liang et al.,2018
Electricity consumption anomaly detection in aluminium extrusion processes
Energy saving
Parameter transfer
Inductive
FCN
Single source
Copiaco et al.,2023
Building’s energy consumption anomaly detection
Energy saving
Parameter transfer
Inductive
MAlexNet-40
Single source
Xu et al.,2021
Power consumption anomaly detection
Energy saving
Mapping transfer
Transductive
Cluster-based DAN
Single source
Simone and Amigoni,2021
Building’s power consumption anomaly detection
Energy saving
Parameter transfer
Inductive
LSTM
Single source
Xiong et al.,2018
Aircraft flight anomaly detection
Infrastructure facilities monitoring
Parameter transfer
Inductive
LSTM
Single source
Pan et al.,2023
Anomaly identification for bridge groups
Infrastructure facilities monitoring
Parameter transfer
Inductive
CNN
Single source
Dhillon and Haque,2020
Network intrusion detection
Infrastructure facilities monitoring
Parameter transfer
Inductive
CNN+LSTM
Single source
Figure 4: Overview of quantity structure of transfer learning problem setting, deep transfer learning approach categories, and label availability in the surveyed industrial domains.
4.2 Manufacturing process monitoring
Manufacturing process monitoring is crucially important to ensure high-quality products and low rejection rates. For example, in injection molding machines, sensors are installed to detect molding conditions in the cavity, such as cavity pressure and temperature. These signals are used to analyze in particular the mold filling and solidification process for each produced part. Such cyclic processing data can also be seen in metal machining (cutting force signal) or joining of parts (joining force signal). Currently, parameter transfer is predominantly used for manufacturing processes (Lockner and Hopmann, 2021; Tercan et al., 2018, 2019; Lockner et al., 2022; Maschler et al., 2021; Abdallah et al., 2021, 2023; Hsieh et al., 2019; Maschler et al., 2021).
In injection molding, parameter transfer is applied to transfer the knowledge from one or more source domains to solve tasks in a target domain (Lockner and Hopmann, 2021; Tercan et al., 2019; Lockner et al., 2022). Similarly, Tercan et al. (2018) build a bridge between simulated data and real data using parameter transfer in injection molding. Maschler et al. (2021) compare different DNNs for anomaly detection tasks on metal forming datasets. Further, they propose a deep transfer learning framework aiming to transfer knowledge between tasks. However, the proposed architecture is not validated. Later, Maschler et al. (2021) apply continuous learning on the same dataset by transferring knowledge from several source tasks to a target task to train a deep learning algorithm capable of solving both source and target tasks. Abdallah et al. (2021, 2023) apply parameter transfer to monitor operation status of manufacturing testbeds with vibration sensor data. Hsieh et al. (2019) transfer knowledge across three chambers in a production line to detect anomalous time series data. Results show reduced training time and improved detection accuracy through transfer learning.
4.3 Predictive Maintenance
Predictive maintenance aims to predict the necessity of maintenance before production is negatively impacted by a failure. Tasks involve monitoring equipment to anticipate maintenance requirements (i.e., predict likely future failure) to optimize maintenance schedules (Serradilla et al., 2022). Time series anomaly detection is often used in respective systems to identify abnormal patterns or behaviors in operation that may indicate the need for maintenance, such as increasing noise, vibrations, etc.
Mao et al. (2020) use mapping transfer with a Sparse Auto-Encoder (SAE) for motor vibration anomaly detection. A transformation from the source and target space to a common latent feature space is learned by MMD to make the feature distribution of two domains as identical as possible. Similarly, Wen et al. (2019) also used mapping transfer with an SAE architecture for fault detection of rotation bearings, using an MMD regularizer to extract a common feature representation. Subsequently, they propose a new MU-Net architecture to deal with multivariate time series anomaly detection tasks (Wen and Keyes, 2019). First, they pre-train a U-Net (Ronneberger et al., 2015) on a large time series dataset for an anomaly detection task. Then, they propose a new model MU-Net, built upon U-Net, wherein each channel they can use the pre-trained U-Net through fine-tuning to transfer knowledge for multivariate time series anomaly detection.
In another application, parameter transfer is used to predict the remaining useful life for tools in manufacturing (Sun et al., 2019). An SAE network is first trained to predict the remaining useful life of a cutting tool on retrospectively acquired data in an offline process. The trained network is then transferred to production with a new tool in operation for online remaining useful life prediction. A 2D CNN-LSTM (Hochreiter and Schmidhuber, 1997) hybrid architecture for fault detection is presented. The model is trained on a fault dataset, and then parameter transfer is applied to target datasets with a different set of conditions. The result shows that transfer learning based hybrid deep learning significantly reduces the training time and is highly suitable for real-time industrial fault diagnosis in various environments. Similarly, parameter transfer is implemented to reduce the gap between different industrial environments (Zabin et al., 2023; Xu et al., 2019). Xu et al. (2019) use a stacked SAE to extract general features from source data and a digital-twin-assisted fault diagnosis approach is presented to transfer knowledge from virtual space to physical space for real-time use. Here, a DNN model is first fully trained in virtual space and then migrated to the physical space using deep transfer learning for real-time use.
The surveyed literature proves that deep transfer learning is a research field that could simplify the life cycle of predictive maintenance systems and facilitate DNN model reusability by reducing the required data and training time, helping adapt them to solve similar tasks.
4.4 Energy management
Energy management deals with systems that detect abnormal excessive consumption caused by end-users’ unusual behavior or malfunction of faulty devices or systems (Copiaco et al., 2023). The goal is to develop automatic, quick-responding, accurate, and reliable fault detection to save energy and build environmentally friendly systems. Energy anomaly detection systems monitor data during energy generation, transmission, and utilization, in order to ensure normal energy consumption. Xu et al. (2021) design a cluster-based deep adaptation layer to improve a deep adaptation network, effectively reducing the mismatch in transfer learning of spinning power consumption anomaly detection. Liang et al. (2018) successfully build an electricity consumption time series anomaly detection method in aluminum extrusion. Parameter transfer is applied to transfer domain knowledge from another data-sufficient domain. They also find it unnecessary when the target data is sufficient because transferring knowledge decreases prediction accuracy.
4.5 Infrastructure facilities monitoring
Infrastructure facilities monitoring refers to monitoring and maintaining the conditions of infrastructure facilities, such as bridges, buildings, and networks. This can include detecting potential issues or failures. The goal is to minimize the impact of failures on the public or the environment. Pan et al. (2023) apply parameter transfer to make full use of the similarity of the anomalous patterns across different bridges and transfer the knowledge obtained by a CNN model to a small part of target data, achieving high accuracy anomaly detection across bridges. Dhillon and Haque (2020) present a parameter transfer approach towards building a network intrusion detection system based on CNN and LSTM.
5 Discussion
5.1 Challenges
Label availability
Deep transfer learning is built upon deep learning, which usually requires a large amount of labeled data, the more data a model has available for learning, the better it can generalize to new examples. In real-world industrial time series anomaly detection tasks, collecting data is probably easy, but collecting labels is much more expensive and time-consuming, sometimes prohibitively so, leading to the unavailability of sufficient labeled data. Self-supervised learning can be used to re-label a large amount of unlabeled data and thus facilitate transfer learning process. Thus, anomaly detection models usually need to learn in an unsupervised or semi-supervised mode (Goldstein and Uchida, 2016).
Deep learning for imbalanced data
Even if the labels can be collected, anomalies can be extremely rare by design, which poses the risk of training with extremely imbalanced data. A practical problem for anomaly detection in industry is the extremely imbalanced data distribution, in which normal samples dominate in data and abnormal samples only share a small percentage in the whole dataset. Prior research has proven that the effect of class imbalance on classification performance by using deep learning is detrimental (Buda et al., 2018). However, most research studies still ignore such problems, which can result in poor performance regarding the minority class, i.e., abnormal data are misclassified as normal.
Missing relevant data
Another problem is missing relevant data, i.e, some information that has a significant effect on the process from case to case is not even recorded or is too complex to record (i.e. part geometry, machine geometry, or environmental conditions in injection molding processes).
Domain shift
Domain shifts lie at the heart of the deep transfer learning problem, but the dynamic changes in many industrial processes, up to an apparent dissimilarity of source and target data, make the transfer learning task particularly challenging.
Effectiveness of deep transfer learning
The general effectiveness of deep transfer learning is limited by the difficulty of determining which knowledge or to what extent the knowledge should be transferred from source to target task. Unlike natural language processing, pretraining a language model on a large corpus of text data can help the model learn the statistical patterns and semantic and syntactic representations of words and sentences, which can be used for new natural language processing tasks with a few data. For industrial time series, due to data privacy, large available public datasets usually do not exist, or they cannot be used even because of a large domain gap between different datasets and tasks. In this case, transferring all of the knowledge may not be beneficial, as it may be irrelevant. In the worst case, this can lead to negative transfer (Pan and Yang, 2010; Smith-Jentsch et al., 2001), in which the extracted knowledge harms the new task-learning. This requires assessing how source and target tasks are related, carefully selecting the knowledge to be transferred, and selecting the proper means to implement this transfer.
5.2 Directions for anomaly detection solution design
Data preprocessing
How data preprocessing should be conducted is an open question. For industrial applications, some researchers consider directly using time series data as input for training to be inefficient, thus suggesting deriving or selecting features from time series data by statistical methods or human experience to decrease the complexity of the dataset dramatically. On the other hand, this crops a lot of potentially useful information, e.g., the time series trend. Some researchers use machine parameters as features of the manufacturing process instead of using process data collected by sensors (Lockner and Hopmann, 2021; Tercan et al., 2018, 2019; Lockner et al., 2022). Others try different transformations of raw time series data, a common way being to transform 1D time series data to 2D image data (Liao et al., 2021; Wang et al., 2021; Zabin et al., 2023) or transforming time domain signals otherwise into the frequency domain (Lukic et al., 2016). However, as large-scale computation power and storage become cheaper and more accessible, it is becoming increasingly common to use deep learning techniques to directly process time series data (Xu et al., 2019; Maschler et al., 2021).
Data augmentation
Data augmentation is useful for deep learning models because it can help to prevent overfitting. For deep transfer learning, when a model becomes too closely adapted to the specifics of the source domain, it may not be able to generalize well to some examples in the task domain. One important technique is to acquire effective synthetic data, e.g., using a simulation process or model to explore potential anomalous conditions by simulating industrial processes under parameters that cannot yet be experienced in the real-world. High fidelity and reliability of simulation data can provide training data at low cost and mitigate the problem of insufficient samples for deep transfer learning (Xu et al., 2019). Another way to generate effective synthetic data is to use GANs. GANs are trained on normal data only to generate indistinguishable samples from which abnormal samples are distinguished during the testing stage of the overall anomaly detection system based on their deviating data distribution (Jiang et al., 2019). To increase the number of anomalous samples and thus the robustness of the anomaly detection model, the technique of adversarial perturbation known from computer vision (Goodfellow et al., 2015) can be used.
Data imbalance
DNNs perform well when they are trained on balanced datasets. However, in practice, it is difficult to get sufficient anomalous data for anomaly detection tasks. For example, manufacturing process is usually in a healthy state due to the pre-designed and optimized operation. Several ways exist to address the imbalanced dataset for time series anomaly detection. To deal with data imbalance, one way is to oversample the minority class, i.e., to randomly replicate samples from the minority class to equalize the number of samples from each class in each batch. Synthetic Minority Over-sampling Technique is an advanced method that creates synthetic samples to force the decision region of the minority class to become more general (Chawla et al., 2002). This technique is widely used in anomaly detection tasks in industry (Ijaz et al., 2018; Mokhtari et al., 2021). Apart from oversampling, the resampling strategy is also frequently used to assign a higher probability to abnormal samples and evenly select the same amount of samples from both classes in each batch. Finally, a weighted loss can be used that balances the loss for the abnormal and normal class in supervised anomaly detection (Buda et al., 2018).
5.3 Directions for deep transfer learning implementation
When shall deep transfer learning be used?
(1) Limited data availability: If the amount of data available for a specific task is limited, pre-training on related source data can learn general features that can be transferred to the specific learning task in the target domain. (2) Similar domains: deep transfer learning is well suited for similar source and target domains. (3) Limited resources (time and compute): Using parameter transfer here is recommended if pre-trained models exist.
When not to use deep transfer learning?
(1) Different tasks: If the target learning task is vastly different from the source learning task, deep transfer learning may not be appropriate. For example, if one wants to train a model for natural language processing on a new dataset, using a pre-trained model that has been trained on image data will not be useful. (2) High domain shift: If there is a large difference between the source and the target domain, deep transfer learning may not be effective. This can happen when the data distributions, features, or labels are vastly different. (3) Abundance of labeled data: If there are enough data for the new task, it may be more effective to train a model from scratch (Liang et al., 2018).
What model architecture to choose?
We suggest selecting the model architecture mainly depending on the data size and label availability, starting from a relatively small network and moving gradually to more complex DNNs. It is important to effectively capture the temporal dependencies and extract respective features of time series data. LSTMs are used heavily for detecting temporal dependencies in time series data (Zabin et al., 2023; Panjapornpon et al., 2023; Xiong et al., 2018). CNNs are also effective in extracting time series features (Yao et al., 2022; Zgraggen et al., 2021). For semi-supervised settings, CNN-based auto-encoders are trained to reconstruct the original data (Serradilla et al., 2021).
Another aspect of deep transfer learning implementations is often the limited computing power of hardware platforms, such as embedded systems in industrial applications. Sensor data are typically acquired using resource-constrained edge processing devices that struggle with computationally intensive tasks, especially when updating a DNN model. One possible solution is federated learning, perhaps the most popular framework, mainly due to its feature of leveraging data while still preserving their privacy (Wang et al., 2018; Amiri and Gündüz, 2019). The technology enables a more collaborative approach to ML while preserving user privacy by storing data decentralized on distributed devices rather than on a central server. Combining deep transfer learning with federated learning is a promising and powerful combination in the abovementioned industrial applications.
Beyond transfer learning
Foundation models such as transformers (Vaswani et al., 2017), diffusion models (Rombach et al., 2022) and SAM (Kirillov et al., 2023) demonstrate emerging properties such as in-context learning (Brown et al., 2020) and complex cross-modality conditioning. This is achieved by training complex and often auto-regressive models with massive amounts of data, although the precise mechanisms that lead to this are not well understood. Some of those models generalize to new settings and tasks, without an explicit element of transfer learning. Thus, the application of foundational models in industrial time series analysis has the potential to reduce and eventually eliminate the need to explicitly account for changes in the domain in modeling, and instead, the foundational models will provide the transfer capability. To not only detect anomalies but also identify failure modes and elicit an appropriate intervention, AI systems must have some form of understanding, a world model, or, in other words, the AI has to implicitly or explicitly model causal relations. Counterfactual inference incorporates causal relations between observations and interventions which allows making predictions of outcomes that were never seen during training (Vlontzos et al., 2023).
6 Conclusions
In this survey, we presented a comprehensive overview of (deep) transfer learning by defining transfer learning problem settings and categorizing the state-of-the-art deep transfer learning approaches. Equipped with this foundation, we selected representative examples of the landscape of fielded applications to provide practitioners with a guide to the field and possibilities of industrial time series anomaly detection. After carefully discussing open challenges, we gave practical directions for time series anomaly detection solution design and deep transfer learning implementation. We found that current applications focus on simple cases with simple datasets, neural network structures, and deep transfer learning schemes. Despite this, the survey suggests that deep transfer learning approaches have huge potential and promise for solving more complex and dynamic anomaly detection tasks in the industry. As the field is still in an early stage, more R&D is expected to fully realize the potential of deep transfer learning in increasingly complex settings. Future work should focus on developing robust transfer learning schemas and methods that can handle more complex and dynamic tasks. The following directions hold the greatest potential for future work:
Automatic selection of transferable features
(Long et al., 2015) It refers to methods for selecting and transferring only the relevant knowledge from the base model. This could involve the use of techniques such as selective fine-tuning and distillation to identify the most important features and knowledge learned from source domains (Yosinski et al., 2014; Ge and Yu, 2017).
Investing into more complex deep transfer learning schemes and DNN architectures
Most deep transfer learning approaches applied in industry focus on the parameter transfer approach as it is conceptionally the simplest and readily applicable by interdisciplinary teams without ML research experience. It seems promising to invest in testing more appropriate deep transfer learning approaches according to different use cases, such as mapping transfer, adversarial transfer, etc. The same applies to testing diverse DNN architectures besides straightforward ones.
Data-centric approach to real-time anomaly detection
The data-centric approach focuses on improving ML models by ensuring high-quality labeled data (Stadelmann et al., 2022) using techniques such as re-labeling, re-weighting, or data augmentation (Luley et al., 2023). Currently, a human-in-the-loop solution is still needed, frameworks have been proposed to assist annotators with graph-based algorithms such as nearest neighbor graphs (Bai et al., 2021), decision trees (Liu et al., 2021), or factor graphs (Kang et al., 2022). Although these methods have proven to be effective, a more automated process is a goal for future research.
Integration with other ML methods
To build robust AI approaches to solve time series anomaly detection in industry, only focusing on transfer learning will not be sufficient. Combinations with other ML approaches are needed in the future, such as continuous learning, meta-learning, and federated learning.
Acknowledgments
We would like to acknowledge Claudio Riginio for his constructive comments on an earlier draft of this paper and for his assistance with the illustrations. This work has been supported by Innosuisse grant 62174.1 IP-ENG “DISTRAL”.

### v2

@@
 Standard deep learning methods are suitable to solve a specific task given a specific type of data.
-During training, the algorithms demand large volumes of labeled training data.
-However, due to the dynamic nature of processes and the environment, it is impractical to acquire the needed data for standard deep learning training for every slightly different case anew.
+During training, deep learning demands large volumes of labeled data.
+However, due to the dynamic nature of the industrial processes and environment, it is impractical to acquire large-scale labeled data for standard deep learning training for every slightly different case anew.
 Deep transfer learning offers a solution to this problem.
-By leveraging knowledge from related tasks and accounting for variations in data distributions, this learning framework solves new tasks even with little or no additional labeled data.
+By leveraging knowledge from related tasks and accounting for variations in data distributions, the transfer learning framework solves new tasks with little or even no additional labeled data.
 The approach bypasses the need to retrain a model from scratch for every new setup and dramatically reduces the labeled data requirement.
-This survey provides an in-depth review of deep transfer learning, examining the problem settings of transfer learning and classifying the prevailing deep transfer learning methods.
-Moreover, we delve into applying deep transfer learning in the context of a broad spectrum of time series anomaly detection tasks prevalent in primary industrial domains, e.g., manufacturing process monitoring, predictive maintenance, energy management, and infrastructure facility monitoring.
-We conclude this survey by underlining the challenges and limitations of deep transfer learning in industrial contexts.
-We also provide practical directions for solution design and implementation for these tasks, leading to specific, actionable suggestions.
-Keywords deep transfer learning  $\cdot$ time series analysis  $\cdot$ anomaly detection  $\cdot$ manufacturing process monitoring  $\cdot$ predictive maintenance
+This survey first provides an in-depth review of deep transfer learning, examining the problem settings of transfer learning and classifying the prevailing deep transfer learning methods.
+Moreover, we delve into applications of deep transfer learning in the context of a broad spectrum of time series anomaly detection tasks prevalent in primary industrial domains, e.g., manufacturing process monitoring, predictive maintenance, energy management, and infrastructure facility monitoring.
+We discuss the challenges and limitations of deep transfer learning in industrial contexts and conclude the survey with practical directions and actionable suggestions to address the need to leverage diverse time series data for anomaly detection in an increasingly dynamic production environment.
 1 Introduction
-The fourth industrial revolution — Industry 4.0 (Kagermann et al., 2011), that is characterized by increasing efficiency through the digitization of production, automation, and horizontal integration across companies (Roblek et al., 2016), and the advent of connected cyber-physical systems – referred to as internet of things (Wang et al.
-2015; Jeschke et al.
-2017; Dalenogare et al.
-2018), increases the need for autonomous and intelligent process monitoring.
-This can be exemplified by the use case of a smart factory in which industrial processes are transformed to be more flexible, intelligent, and dynamic (Kagermann, 2017), or the use case of decentralized energy production with wind and solar (Abir et al., 2021).
+1.1 motivation and contribution
+The fourth industrial revolution – Industry 4.0 [1], that is characterized by increasing efficiency through the digitization of production, automation, and horizontal integration across companies [2], and the advent of connected cyber-physical systems – referred to as internet of things [3, 4, 5], increases the need for autonomous and intelligent process monitoring.
+This can be exemplified by the use case of a smart factory in which industrial processes are transformed to be more flexible, intelligent, and dynamic [6], or the use case of decentralized energy production with wind and solar [7].
 In these examples, AI-powered anomaly detection integrates the analysis of time series data to detect unusual patterns in the recorded data.
-By identifying parameters that fall outside a window of normal operation, operators can trigger interventions and adjustments to ensure high product quality and safe operations.
-To achieve this, physical properties such as pressure or temperature are monitored and analyzed in real-time.
-Changes in these variables capture drifting and abrupt faults caused by process failures or malfunctions (Park et al., 2020).
+To achieve this, a deep learning architecture is modeled to capture indicators of normal and abnormal operation.
+The learning process involves the analysis of historic time series sensor data of normal and possibly abnormal operations.
+This data is for example used for representation- or reconstruction-based learning.
+After training, the deep learning model represents or reconstructs normal data in a certain way.
+The model is designed in a way that abnormal data–because it is different–is either represented differently from the normal data or reconstructed poorly and thus recognized as an anomaly.
+By identifying operational parameters that fall outside a window of normal interval, operators can trigger interventions and adjustments to ensure high product quality and safe operations.
+To achieve this, physical properties such as pressure or temperature are monitored and analyzed in real-time applications.
+Changes in these variables capture drifting and abrupt faults caused by process failures or malfunctions [8].
 The production process must adapt quickly to changes in production and the environment to meet the requirements for flexibility and dynamics.
-Further use cases exist in such diverse areas as manufacturing monitoring including automatic quality control, predictive maintenance of goods and services, infrastructure monitoring of e.g.
-building energy systems or power plants, digital agriculture, petrochemical process optimization, computer network intrusion detection, or aircraft flight monitoring, to name a few.
-Artificial Intelligence, in particular deep learning, provides competent frameworks to automate intelligent monitoring in order to provide valuable assistance to operators and high-level control systems.
-Leveraging the power of deep learning, formative features of the data – technically referred to as representations (Bengio et al., 2013) – can be captured in a machine-learned model and thereby enable a detailed understanding of variations in standard operations.
-However, in non-trivial and non-stationary conditions, the task or the underlying data may change.
-For example, the monitoring system of a milling machine may in one instance be tasked to detect a blunt tool based on the vibration and in another instance – using the same vibration measurements – to detect insufficient cooling lubricant.
-Knowledge acquired to solve one task in one setting with a given tool, machined part, and type of machine may be transferred to solve the same or similar task in a setting with a different tool, machined part, or type of machine.
-Slowly changing conditions (drifts), abrupt mode changes (for instance due to tool change), and new tasks (such as the detection of another failure mode) may require adjustments to the machine-learned model.
-In these cases, it is desirable to adjust the analysis model without retraining from scratch, as it is costly or impractical to acquire sufficient training data to learn the full manifold (Maschler et al., 2022).
-Transfer learning is a machine learning framework to achieve this (Maschler and Weyrich, 2021; Zhuang et al., 2021; Pan and Yang, 2010; Tan et al., 2018; Yosinski et al., 2014).
-As depicted in Figure 1, data and algorithms from a related task may be leveraged in a new one.
-By accounting for changes in data distributions and tasks, or leveraging existing models, knowledge learned from related tasks can be used to improve performance on new tasks instead of retraining a model for each individual application from scratch.
+Further use cases exist in a wide range of diverse fields, such as manufacturing monitoring including automatic quality control [9, 10], predictive maintenance of goods and services [11, 12, 13, 14, 15, 16], infrastructure monitoring of building systems [17, 18] and power plant [19], digital agriculture [20], petrochemical process optimization [21], computer network intrusion detection [22], or aircraft flight monitoring [23], to name a few.
+Artificial Intelligence, particularly deep learning, provides competent frameworks with underlying deep neural networks to automate intelligent monitoring and provide valuable assistance to operators and high-level control systems.
+Leveraging the power of deep learning, informative features of the data – technically referred to as representations [24] – can be captured in a machine-learned model and thereby enable a detailed understanding of variations in standard operations.
+However, the task or underlying data may change under non-trivial and non-stationary conditions.
+For instance, the monitoring system of a milling machine may be assigned the task of identifying a blunt tool based on vibration in one scenario, and in a different scenario, it may utilize the same vibration measurements to detect insufficient cooling lubricant.
+Knowledge acquired to solve one task in one setting with a given tool, machined part, and type of machine may be transferred to solve the same or similar task in another setting with a different tool, machined part, or type of machine.
+Slowly changing conditions (drifts), abrupt mode changes (for instance, due to tool change), and new tasks (such as the detection of another failure mode) may require adjustments to the deep learning model.
+In these cases, it is desirable to adjust the analysis model without retraining from scratch, as it is costly or impractical to acquire sufficient training data to learn the full manifold [25].
+Transfer learning is a machine learning framework to achieve this [26, 27, 28, 29, 30].
+As depicted in Fig.
+1, data and algorithms from one task may be leveraged in a new related one.
+By accounting for changes in data distributions and tasks or leveraging existing models, knowledge learned from related tasks can be used to improve performance on new tasks instead of retraining a model for each individual application from scratch.
 This transfer-learning-boosted modeling forms the basis for identifying anomalies that deviate from established patterns in a non-trivial manner without full re-training.
@@
 In the production of gear A, a lot of data is available to train a deep learning model that helps improve production.
-In the production of gear B, data availability is more limited, and the traditionally trained deep learning model fails to improve production.
-With suitable methods, however, data and algorithms acquired during production of gear A can be leveraged to support improving the production of gear B, because the data and tasks in the production of both gears are related.
-Deep transfer learning (Tan et al., 2018; Yu et al., 2022) extends the transfer learning paradigm by leveraging deep learning models.
+In the production of gear B, data is more limited, and the traditionally trained deep learning model fails to improve production.
+With suitable transfer learning methods, however, data and algorithms acquired during the production of gear A can be leveraged to support improving the production of gear B because the data and tasks in the production of both gears are related.
+Deep transfer learning [29, 31] extends the transfer learning paradigm by leveraging deep learning.
 In industrial contexts, it ensures optimal production even as production conditions shift.
 This dynamic adaptability is key in maintaining the effectiveness of anomaly detection systems in the dynamic environment that characterizes industrial applications including the broad categories of manufacturing process monitoring, predictive maintenance, energy management, and infrastructure facility monitoring as detailed in Section 4.
-In this survey, we review the foundations of deep transfer learning to equip the reader with working knowledge of the main principles and intuition for ideas.
-Further, we provide a comprehensive overview of the current state of the art of deep transfer learning approaches for time series anomaly detection for industrial applications.
-Our main contribution is a systematic review of research work on real-world industrial applications.
-For these, we discuss potential, challenges, and limitations and give directions for future work and potential.
-The paper is organized as follows: First, we introduce a taxonomy of transfer learning problem settings and further categorization of deep transfer learning approaches (Section 2).
+This survey is a non-systematic yet application-oriented review with a narrow focus on deep transfer learning for anomaly detection in time series in the industry.
+Our main contributions are as follows:
+  • We categorize transfer learning problem settings and then systematically summarize deep transfer learning approaches into four categories.
+With the foundations of deep transfer learning, we equip the reader with a working knowledge of the main principles and intuitions.
+  • We analyze the recent literature and provide a comprehensive overview of the current state of the art of deep transfer learning approaches for time series anomaly detection for main industrial applications.
+  • We discuss potential challenges and limitations and then give directions for future work with actionable recommendations for AI practitioners and decision-makers.
+To our knowledge, this is the first survey of deep transfer learning in the narrow context of industrial time series anomaly detection.
+The review describes the underlying methodological principles and methods within a generic taxonomy and discusses practical implications for AI practitioners to make informed decisions.
+We cover multiple areas of application, including manufacturing monitoring, maintenance prediction, and infrastructure monitoring.
+The rest of the paper is organized as follows.
+First, we provide an overview of transfer learning by introducing a taxonomy of transfer learning problem settings and further categorizing deep transfer learning approaches (Section 2).
 Then, we describe the task of anomaly detection in time series (Section 3) in selected industrial applications (Section 4).
 To conclude, we discuss current challenges, limitations, and future research directions (Sections 56) in the field.
+1.2 Survey methodology
+We seek to identify application-oriented peer-reviewed literature in the intersection of transfer learning as the learning framework, time series as the data domain, and anomaly detection as the task (Fig.
+2).
+To execute the selection process of literature, we search related terms on Google Scholar, Scopus, Elsevier, and IEEE databases.
+Based on the title, we pick those papers that may fit the narrow topic into a pre-selection list.
+Eventually, we included publications matching the topic according to the abstract and screening of the content.
+Along the reviewed topical papers, we include contextually relevant papers such as deep learning approaches that are agnostic to data types and tasks.
+For deep transfer learning in general, we searched the keywords “transfer learning” and “deep transfer learning”.
+Specifically, we focus more on deep transfer learning approaches.
+Then, we switch to the application-oriented cases where deep transfer learning is applied to tackle time series anomaly detection in the main industrial applications.
+To achieve this, we search queries like “deep transfer learning for time series anomaly detection” and “deep transfer learning for predictive maintenance”.
+After searching in the database, we carefully check and screen out the most relevant literature based on the following inclusion/exclusion criteria: (1) We only include the applications that utilize deep transfer learning approaches, instead of traditional transfer learning; (2) We only include publications after 2013; (3) We cover all three main topics in Fig.
+2 (highlighted with cycles), but we specifically focus on the intersection of the three aforementioned domains.
+After carefully screening out, we select 45 papers for deep transfer learning in general and 37 papers for deep transfer learning for anomaly detection in industrial time series.
+Fig.
+3 illustrates the taxonomy in this survey to categorize reviewed studies based on different aspects, including deep transfer learning, time series anomaly detection, industrial applications, current challenges, and future directions.
+Figure 2: Venn diagram of this survey’s focus on the intersection of transfer learning, anomaly detection, and time series analysis.
+Figure 3: A generic taxonomy in this paper to analyze deep transfer learning for industrial time series anomaly detection.
 2 Deep transfer learning
-Transfer learning in a deep learning setting aims to increase the efficiency, performance, and generalization of deep learning models by transferring knowledge from one data set and task to a new one.
-This eliminates the need to train a deep learning model from scratch, which in turn reduces the amount of necessary data and compute required to solve a new task or new data domain.
-In either case, knowledge is transferred from a source to a target domain, as defined below.
-The transfer learning problem settings can be categorized as inductive or transductive transfer depending on the data and task conditions, while we categorize deep learning-based transfer learning approaches into instance transfer, parameter transfer, mapping transfer and domain-adversarial transfer.
-We illustrate them by using two intuitive examples in Figure 2, with more details being elaborated in the following sections.
+2.1 Overview of the field
+[…]
+[…]
 Figure 4: Taxonomy of transfer learning problem settings (left; see Section 2.2 for the definition of terms) and corresponding examples using deep transfer learning approaches (right).
-On the left, we classify transfer learning problems as inductive or transductive transfer settings, mainly depending on the label availability and task similarities.
-Correspondingly, we provide two examples using deep transfer learning methods: In the inductive transfer setting, we collect time series data from screw production and wrench production.
-Labeled screw data (A1) is used to detect collective anomalies (a set of data points behaving differently compared to the entire time series (Choi et al., 2021; Chalapathy and Chawla, 2019), further explained in Section 3).
+On the left, we classify transfer learning problems as inductive or transductive transfer settings Correspondingly, we provide two examples using deep transfer learning methods: In the inductive transfer setting, we collect time series data from screw production and wrench production.
+Labeled screw data (A1) is used to detect collective anomalies (a set of data points behaving differently compared to the entire time series [32, 33], further explained in Section 3).
 Then, parameter transfer (Section 2.3.2) is applied to transfer knowledge by fine-tuning the pre-trained model from labeled screw data to detect point anomalies (further explained in Section 3) on labeled wrench data (A2).
@@
 However, the data in B2 significantly differs in appearance from the data in B1.
-To address this issue, instance transfer (further explained in Section 2.2.1) is used.
-Through this learning process, the data in B2 is transformed in a way that makes it compatible with the model which has been trained exclusively on data from B1.
+To address this problem, instance transfer (further explained in Section 2.3.1) is used.
+Through this learning process, the data in B2 is transformed in a way that makes it compatible with the model that has been trained exclusively on data from B1.
 Transfer learning, in this case, is thus achieved by adapting the data to fit the model through domain adaptation rather than adjusting the model to fit new data.
-2.1 Transfer learning problem definition
+2.2 Formal description of deep transfer learning
 Domain $\mathcal{D}$ includes the domain feature space $\mathcal{X}$ and marginal data distribution $P(X)$ as $\mathcal{D}=\{\mathcal{X},P(X)\}$, where $X$ is the domain data, $X=\left\{x_{1},\ldots,x_{n}\right\}\in\mathcal{X}$.
@@
 The function $f_{\mathcal{T}}(\cdot)$ can be used to predict the corresponding label of a new instance $x_{i}$.
-The objective predictive function $f_{\mathcal{T}}(\cdot)$ can be learned from domain data and can be interpreted as a form of conditional probability.
+The objective predictive function $f_{\mathcal{T}}(\cdot)$ learned from domain data can be interpreted as a form of conditional probability.
 Thus, the learning task can be rewritten as $\mathcal{T}=\{\mathcal{Y},P(Y|X)\}$, where $P(Y|X)$ is used as a likelihood measure to determine how well a given data set $X$ fits with a corresponding class label set $Y$.
-We largely follow the definition of transfer learning by Pan and Yang (2010) and Zhuang et al.
-(2021).
+In the surveyed literature on transfer learning for anomaly detection in industrial applications, transfer learning methods from other fields, such as computer vision and natural language processing, were adopted.
+We therefore use a generic classification scheme for transfer learning methods.
+We largely follow the definition of transfer learning in literature [27, 28].
 Given a source domain $\mathcal{D}_{S}$ and learning task $\mathcal{T}_{S}$, as well as a target domain $\mathcal{D}_{T}$ and learning task $\mathcal{T}_{T}$, transfer learning aims to improve the performance of the predictive function $f_{\mathcal{T}}(\cdot)$ in $\mathcal{D}_{T}$ by transferring knowledge from $\mathcal{D}_{S}$ and $\mathcal{T}_{S}$, where $\mathcal{D}_{S}\neq\mathcal{D}_{T}$ and/or $\mathcal{T}_{S}\neq\mathcal{T}_{T}$.
 Usually, the size of source dataset is much smaller than target dataset.
-This definition of transfer learning can be broadened, i.e., the target task can profit from multiple source domains.
+This definition of transfer learning can be broadened, i.e., the target task can benefit from multiple source domains.
 Transfer learning is thus the idea of making the best use of related source domains to solve new tasks.
 In contrast, traditional machine learning (ML) methods learn each task separately from scratch, and each respective model can only be applied to the corresponding task.
-We define a taxonomy of transfer learning problems settings as shown in Figure 2 mainly depending on the label availability in the two domains to be easily applicable to the requirements of a case at hand (compare different definitions for other purposes in the literature Pan and Yang (2010); Tan et al.
-(2018); Zhuang et al.
-(2021)).
-We differentiate it into inductive and transductive transfer learning (Pan and Yang, 2010) [^1] In this survey, we do not consider unsupervised learning scenarios since either source labels or target labels are provided for most industrial applications..
+We define a taxonomy of transfer learning problem settings as shown in Fig.
+4 mainly depending on the label availability in the two domains to be easily applicable to the requirements of a case at hand (compare different definitions for other purposes in the literature [27][29]).
+We differentiate it into inductive and transductive transfer learning [28].
 Inductive transfer learning is applied when the target task is different from the source task, i.e., $\mathcal{T}_{S}\neq\mathcal{T}_{T}$ (meaning that $\{\mathcal{Y_{S}}\neq\mathcal{Y_{T}}\}$ or $\{P(Y_{S}|X_{S})\neq P(Y_{T}|X_{T})\}$).
-The conditional probability distribution is induced with labeled training data in the target domain (Torrey and Shavlik, 2009).
-A corresponding example is illustrated as Scenario A in Figure 2, where the learning tasks are different and the goal of transfer learning is to recognize point anomaly from the collective anomaly task.
-Related areas of inductive transfer learning are multi-task learning (Caruana, 1997; Ruder, 2017) and sequential learning, depending on whether tasks are learned simultaneously or sequentially.
+The conditional probability distribution is induced with labeled training data in the target domain [34].
+A corresponding example is illustrated as Scenario A in Fig.
+4, where the learning tasks are different and the goal of transfer learning is to recognize point anomaly from the collective anomaly task.
+Related areas of inductive transfer learning are multi-task learning [35, 36] and sequential learning, depending on whether tasks are learned simultaneously or sequentially.
 Transductive transfer learning is applied when the source and target tasks are the same, while the source and target domain are different, i.e., $\mathcal{T}_{S}=\mathcal{T}_{T}$ and $\mathcal{D}_{S}\neq\mathcal{D}_{T}$ (meaning that $\{\mathcal{X_{S}}\neq\mathcal{X_{T}}\}$ or $\{P(X_{S})\neq P(X_{T})\}$.
-A subcategory is domain adaptation (Kouw and Loog, 2019) when the feature space of source and target data are the same but the corresponding marginal distributions are different (i.e., $\{\mathcal{X_{S}=X_{T}}\}$ and $\{P(X_{S})\neq P(X_{T})\}$).
-Scenario B in Figure 2 is an example of transductive transfer learning where the learning tasks are identical, and the goal of transfer learning is to recognize contextual anomalies in an unlabelled data set.
-Other learning paradigms closely related to transfer learning are listed below:
-Multi-task learning is a machine learning technique where a single model is trained on multiple tasks simultaneously.
-The idea is to improve the performance of the model by learning a shared representation that captures the features between all tasks.
-Continuous learning (Parisi et al., 2019) is a learning process where the model continuously learns new tasks from previous tasks over time without forgetting how to solve previous tasks.
-To some extent, continuous learning can be seen as a sequential transfer learning process, with the constraint to preserve the performance on the previous tasks which leads to an accumulation of knowledge over time.
-Few-shot learning (Wang et al., 2020) is a type of machine learning where a model can learn and perform well on a new task with only a limited number of labeled samples.
-In extreme cases, the model can learn with one label (Fei-Fei et al., 2006) and without any label (Lampert et al., 2009).
-While transfer learning usually involves reusing the model from relevant tasks and continuing training on the target dataset.
-Meta-learning (Finn et al., 2017; Hospedales et al., 2022) is a machine learning technique that focuses on the learning process.
-It is known as “learning to learn”.
-For meta-learning, models are trained on a different set of tasks instead of a set of data in the traditional machine learning setting.
-In this sense, meta-learning can be seen as a form of transfer learning because it involves transferring knowledge from task to task.
-Knowledge distillation (Gou et al., 2021) effectively learns a small model trained to mimic the behavior of a larger, more complex model.
-The knowledge learned by the larger model can be transferred to the smaller model, which can then be used for the target task, e.g., on a less powerful edge device.
-Self-supervised learning (Liu et al., 2023; Bai et al., 2021) involves training a model to predict some aspect of the input data without any external supervision.
-The learned representations can be used for various downstream tasks, including those that involve transferring knowledge from one domain to another.
+A subcategory is domain adaptation [37] when the feature space of source and target data are the same but the corresponding marginal distributions are different (i.e., $\{\mathcal{X_{S}=X_{T}}\}$ and $\{P(X_{S})\neq P(X_{T})\}$).
+Scenario B in Fig.
+4 is an example of transductive transfer learning where the learning tasks are identical, and the goal of transfer learning is to recognize contextual anomalies in an unlabelled data set.
 2.3 Deep transfer learning approaches
-Since deep neural networks (DNNs) can learn useful feature representations from large amounts of data through back-propagation (Bengio et al., 2013), they have been widely adopted for tackling complex problems, which involve large-scale and high-dimensional data, also in practice (Schmidhuber, 2015; Lukic et al., 2016; Stadelmann et al., 2018; Schmidhuber, 2022).
+Since deep neural networks (DNNs) can learn useful feature representations from large amounts of data through back-propagation [24], they have been widely adopted for tackling complex problems in practice [38, 39, 40, 41], which involve large-scale and high-dimensional data.
 Deep transfer learning methods implement transfer learning principles within DNN and, among other things, enable deep learning based analysis pipelines to be applied to new datasets.
-On a high level, deep transfer learning approaches can be divided into data-driven and model-driven ones.
-Data-driven approaches focus on transferring knowledge by transforming and adjusting the data instances.
-Model-driven approaches leverage DNNs to develop domain-invariant features by reducing the feature discrepancy between source and target domain data and then transferring generalized knowledge to new tasks.
-Following the taxonomy of (Tan et al., 2018; Pan and Yang, 2010; Shi et al., 2022), we divide deep transfer learning approaches further into $4$ categories: instance transfer, parameter transfer, mapping transfer, and domain-adversarial transfer, as illustrated in Table 1.
-Instance transfer and mapping transfer are data-driven approaches, parameter transfer is a model-driven approach, while domain-adversarial transfer is a combination of both.
+Based on the transferring techniques in the surveyed literature, we access how knowledge is shared across domains and help increase the performance in the target task or domain.
+we divide deep transfer learning approaches further into $4$ categories: instance transfer, parameter transfer, mapping transfer, and domain-adversarial transfer, as illustrated in Table 1.
+Furthermore, instance transfer, mapping transfer, and domain-adversarial transfer can be described as data-driven approaches.
+They focus on transferring knowledge by leveraging a large amount of data.
+It usually involves transforming and adjusting the data instances or manipulating data from different domains by feature alignment, feature mapping, etc.
+On the other hand, parameter transfer is a model-driven approach, which places more emphasis on understanding the underlying structure and dynamics of the data.
+It usually involves transferring the parameters of pre-trained model from source domain to target domain.
 Table 1: Overview of deep transfer learning approaches with references.
@@
 Instance transfer
-Augmenting target data by transforming data instance from the source domain
-He et al., 2022; Amirian et al., 2021; Wang et al., 2018
+Augmenting target data by transforming data instance from the source domain to the target domain
+[42, 43, 44]
 Parameter transfer
-Transfer learned parameters of a pre-trained model from source domain and adapting the model for target domain
-Bommasani et al.,2022; Devlin et al.,2019; Vaswani et al.,2017; Dou et al.,2022; Alexandr et al.,2021; Yosinski et al.,2014; Guo et al.,2019; Sager et al.,2022; Brown et al.,2020
+Transfering learned parameters of a pre-trained model from source domain and adapting the model for target domain
+[45, 46, 47, 48, 49, 50, 17, 51, 52, 53]
 Mapping transfer
 Reducing feature discrepancies between source and target domains by minimizing the distance between mapped features in the latent space
-Zhuang et al., 2021; Tzeng et al., 2014; Long et al., 2017, 2015; Zhang et al., 2015; Venkateswara et al., 2017
+[27, 54, 55, 56, 57, 58, 59]
 Domain-adversarial transfer
 Extracting an indiscriminative feature representation between source and target domain through adversarial training
-Tzeng et al., 2015; Ganin et al., 2017; Ajakan et al., 2015; Tzeng et al., 2017
+[60, 61, 62, 63, 64, 65]
 2.3.1 Instance transfer
@@
 A typical approach is to re-create some labeled data from the source domain.
-He et al.
-(2022) propose an instance-based deep transfer learning model with attention mechanism to predict stock movement.
-They first create new samples from the source dataset that are similar to the target samples by using attention weights, and then train on the created samples and target training samples for prediction tasks.
-Amirian et al.
-(2021) introduce an innovative instance transfer method for domain adaptation.
+For example, He et al.
+propose an instance-based deep transfer learning model with an attention mechanism to predict stock movement [42].
+They first create new samples from the source dataset that are similar to the target samples by using an attention network and then train another network on the created samples and target training samples for prediction tasks.
+Since two networks are trained separately for different tasks, it needs further investigation to what extent the generated samples can contribute to the prediction task.
+Amirain et al.
+introduce an innovative instance transfer method for domain adaptation [43].
 They propose an effective auto-encoder model with a pseudo-label classifier to reconstruct new data instances that obtain general features across different datasets for medical image analysis.
 Taking another avenue, Wang et al.
-(2018) exclude the source data that have a bad impact on training target data.
-Specifically, they choose a pre-trained model from a source domain, estimate the impact of all training samples in the target domain, and remove samples that lower the performance of the model.
+exclude the source data that negatively impacts training target data.
+Specifically, they choose a pre-trained model from a source domain, estimate the impact of all training samples in the target domain, and remove samples that lower the model’s performance.
+Then, the optimized training data is used for fine-tuning.
+The experiments are conducted on large image datasets [44].
+Instead of transferring the data, the approach excludes certain samples based on the pre-trained model’s predictions.
+Additional validation is required in industrial environments, especially when only a few data are available in some industrial settings.
 2.3.2 Parameter transfer
-Parameter transfer adapts learned parameters of a pre-trained model to a new model.
+Parameter transfer adapts the learned parameters of a pre-trained model to a new model.
 This assumes that DNNs can get similar feature representations from similar domains.
 Thus, through transferring parts of the DNN layers together with pre-trained parameters and/or hyperparameters, the pre-trained model is used as a base model to further train on target domain data and solve different learning tasks.
-Particularly, parameter transfer has gained popularity in computer vision and natural language processing, where large models are pre-trained on large datasets (Bommasani et al., 2022).
-In natural language processing, BERT (Devlin et al., 2019) and GPT-3 (Brown et al., 2020) that are based on the transformer architecture (Vaswani et al., 2017) can be fine-tuned for a variety of natural language processing tasks, including content generation (Dou et al., 2022), language translation (Sun et al., 2021), question answering (Glass et al., 2020), and summarization (Alexandr et al., 2021).
-Yosinski et al.
-(2014) investigated the general transferability of DNNs.
-Experiment results show that transferring features from source to target domain leads to improved generalization in networks compared to those trained solely on the target dataset.
+Particularly, parameter transfer has gained popularity in computer vision and natural language processing, where large models are pre-trained on large datasets [45].
+In natural language processing, for example, BERT [46] and GPT-3 [53] are based on the Transformer architecture [48] which can be fine-tuned for a variety of downstream tasks, including content generation [49], language translation [66], question answering [67], and summarization [50].
+In computer vision, Yosinski et al.
+investigate the general transferability of CNNs in image recognition [30].
+They analyze the transferring effect by fine-tuning or freezing a certain amount of layers in the networks.
+Experimental results show that transferring features from source to target domain improves network generalization compared to those trained solely on the target dataset.
+Additionally, they quantify the model performance by assessing how features at what layers transfer from one task to another.
+It is surprising to find that transferring a pre-trained network from any number of layers can produce a boost for fine-tuning on a new dataset.
+However, the experiments are only conducted on certain image datasets, and Tuggener et al.
+show [68] the limits of parameter transfer when the chosen architecture is overfitted on the particularities of certain large-scale datasets.
 Unlike the typical way of fine-tuning a pre-trained model, Guo et al.
-(2019) propose the adaptive fine-tuning approach SpotTune to find the optimal fine-tuning strategy for the target task.
+propose an adaptive fine-tuning approach SpotTune to find the optimal fine-tuning strategy for the target task [51].
 Specifically, a policy network is used to make routing decisions on whether to pass the target instance through the pre-trained model.
@@
 Sager et al.
-(2022) propose an unsupervised domain adaptation for vertebrae detection in 3D CT volumes by transferring knowledge across domains during each batch of the training process.
+propose an unsupervised domain adaptation for vertebrae detection in 3D CT volumes by transferring knowledge across domains during the training process [52].
 2.3.3 Mapping transfer
-Mapping transfer refers to learning a related feature representation for the target domain by feature transformation, which includes feature alignment, feature mapping, and feature encoding (Zhuang et al., 2021).
-The goal is to reduce feature discrepancies between source and target domains by minimizing the distance between the distribution of mapped features in the latent space.
-There are various criteria to measure the distribution difference, including Wasserstein distance (Shen et al., 2018), Kullback-Leibler Divergence (Dai et al., 2007), etc.
-Among them, Maximum Mean Discrepancy (MMD) (Tzeng et al., 2014) is most frequently adopted in mapping transfer from the surveyed papers.
+Mapping transfer refers to learning a related feature representation for the target domain by feature transformation, which includes feature alignment, feature mapping, and feature encoding [27].
+The goal is to reduce feature discrepancies between source and target domains by minimizing the distance between the distribution of latent feature representation.
+There are various criteria to measure the distribution difference, including Wasserstein distance [69], Kullback-Leibler Divergence [70], etc.
+Among them, Maximum Mean Discrepancy (MMD) [55] is most frequently adopted in mapping transfer from the surveyed papers.
 The MMD is calculated as the difference between the mean embeddings of the samples in a reproducing kernel Hilbert space associated with a chosen kernel function.
-Added to the target loss function, it measures the difference between two probability distributions and serves as a powerful tool for comparing the similarity of complex, high-dimensional datasets using a wide variety of kernel functions.
-Some previous work focusing on transferred feature extraction/dimensionality reduction using MMD has been done.
+Added to the target loss function, it serves as a powerful tool for comparing the similarity of complex, high-dimensional datasets using a wide variety of kernel functions.
+Previous work has focused on transferred feature extraction/dimensionality reduction using MMD.
+Wang et al.
+focus more on the subdomain of the same subcategory instead of the alignment of the global distribution between source and target domain [54].
+Specifically, they first use the attention mechanism to extract discriminative features that are most related to the fault signal.
+Then, local MMD is applied to transfer knowledge to adjust the distribution of related subdomains under the same category.
 Long et al.
-(2017) base their Joint Adaptation Network on MMD, in which the joint distributions of multiple domain-specific layers across domains are aligned.
+propose their Joint Adaptation Network [56] based on MMD, in which the joint distributions of multiple domain-specific layers across domains are aligned.
 In addition, an adversarial training version was adopted to make distributions of the source and target domains more distinguishable.
 Similarly, Long et al.
-(2015) adopted multi-layer adaptation and proposed Deep Adaptation Networks (DAN).
-In DAN models, the first three convolutional layers are used to extract general features.
-For the last three layers, multi-kernel MMD is used to bridge the cross-domain discrepancy and learn transferable features.
+adopt multi-layer adaptation and proposed Deep Adaptation Networks (DAN) [57].
+The first three convolutional layers are used in DAN models to extract general features.
+For the last three layers, multi-kernel MMD bridges the cross-domain discrepancy and learns transferable features.
 Zhang et al.
-(2015) also based on MMD and proposed a Deep Transfer Network in which two types of layers are used to obtain domain invariant features across domains.
-The shared feature extraction layers learn a shared feature subspace between the source and the target samples, and the discrimination layer is then used to match conditional distributions by classifier transduction.
+propose a Deep Transfer Network in which two types of layers are used to obtain domain invariant features across domains by adding MMD loss.
+The shared feature extraction layers learn a shared feature subspace between the source and the target samples, and the discrimination layer is then used to match conditional distributions by classifier transduction [58].
 Venkateswara et al.
-(2017) proposed Deep Adaptation Hash network, which is fine-tuned from the VGG-F (Chatfield et al., 2014) network.
-Multi-kernel MMD loss trains the Deep Adaptation Hash to learn feature representations that align the source and target domains.
+propose Deep Adaptation Hash network [59], which is fine-tuned from the VGG-F [71] network.
+Multi-kernel MMD loss is employed to train the Deep Adaptation Hash network to learn feature representations that align the source and target domains.
 2.3.4 Domain-adversarial transfer
-Inspired by Generative Adversarial Networks (GANs) (Goodfellow et al., 2014; Schmidhuber, 2020), the goal here is to extract a transferable feature representation that is indiscriminative between source and target domain through adversarial training.
-Adversarial transfer mainly focuses on domain adaptation problems.
+Inspired by Generative Adversarial Networks (GANs) [72, 73], the goal of domain-adversarial transfer is to extract a transferable feature representation that is indiscriminative between source and target domain through adversarial training.
+Adversarial transfer is primarily concerned with addressing domain adaptation problems.
+Soleimani and Nazerfard utilize the GANs framework to perform cross-subject transfer learning [60].
+The generator is used to generate samples that are similar to the target data.
+Meanwhile, the discriminator distinguishes the fake samples from the target samples.
+The classifier is trained to discriminate the labeled source data and fake samples to learn generalized features invariant to source and target domains.
+It is important to note that in real-world applications, training GANs can be unstable due to mode collapse, especially in the case when the source data and target data are unbalanced, and the generator may fail to generate fake samples that can confuse the discriminator.
 Tzeng et al.
-(2015) adopt a domain confusion loss across the source and target domains to learn a domain invariant representation.
+adopt a domain confusion loss across the source and target domains to learn a domain invariant representation [61].
 Ganin et al.
-(2017) propose a new domain adaptation architecture by adding a domain classifier after feature extraction layers.
+propose a new domain adaptation architecture by adding a domain classifier after feature extraction layers [63].
 A gradient reversal layer is used to ensure the similarity of the feature distributions over source and target domains.
+Similarly, Ozyurt et al.
+develop a novel framework for unsupervised domain adaptation of time series data by using contrastive learning and domain-adversarial transfer learning [62].
+A domain classification loss is applied to extract domain invariant features.
+The drawback is that the experiments are designed in a way that the source and target data sizes are similar, whereas in practice, the target data is usually much fewer than the source data.
 Ajakan et al.
-(2015) propose a domain adversarial DNN in which a domain regressor is applied to learn a domain invariant feature representation.
+propose a domain adversarial DNN in which a domain regressor is applied to learn a domain invariant feature representation [64].
 Tzeng et al.
-(2017) use an unsupervised domain adaptation method that combines adversarial learning with discriminative feature learning.
-3 Time series analysis for industrial processes
-Time series analysis encompasses statistical techniques to analyze and interpret sequential temporal data.
-In the context of industrial processes, time series analysis plays a crucial role in automating monitoring and controlling the efficiency, quality, and performance of these processes.
-Specifically, the analysis of time series data can be used for anomaly detection, forecasting, process control, performance assessment, and maintenance scheduling to increase the efficiency of the process.
-Anomaly detection
-According to Hawkins (1980), an outlier is as an observation that deviates significantly from other observations in a way that it is likely that it was generated by a different mechanism.
+use an unsupervised domain adaptation method that combines adversarial learning with discriminative feature learning [65].
+2.4 Related learning paradigms
+Besides the dedicated transfer learning approaches discussed above, there are methods that represent alternative ways to solve tasks across domains or are complementary to the native transfer learning methods.
+  • Multi-task learning is a machine learning technique where a single model is trained on multiple tasks simultaneously.
+The idea is to improve the performance of the model by learning a shared representation that captures the features between all tasks.
+Because the network learns to solve multiple tasks, it may generalize better to new data and tasks.
+  • Continuous learning [74] is a learning process where the model continuously learns new tasks from previous tasks over time without forgetting how to solve previous tasks.
+To some extent, continuous learning can be seen as a sequential transfer learning process, with the constraint to preserve the performance of the previous tasks, which leads to an accumulation of knowledge over time.
+  • Few-shot learning [75] is a type of machine learning where a model can learn and perform well on a new task with only a limited number of labeled samples.
+In extreme cases, the model can learn with one label [76] and without any label [77].
+Whereas, transfer learning usually involves reusing the model from relevant tasks and continuing training on the target dataset.
+  • Domain generalization[78, 79] focuses on developing a generalized model from one or multiple distinct domains to detect unseen target domain data.
+The main goal is to overcome the domain shift problem.
+Domain generalization and transfer learning are both applied to transfer knowledge from source domain to target domain.
+The major difference between transfer learning and domain generalization lies in the utilization of target domain data.
+Transfer learning leverages knowledge from source domain and target domain.
+In contrast, domain generalization solely learns from source domain, without access to the target data.
+  • Meta-learning [80, 81] is known as “learning to learn”.
+For meta-learning, models are trained on a different set of tasks instead of a set of data in the traditional machine learning setting.
+In this sense, meta-learning can be seen as a form of transfer learning because it involves transferring knowledge from task to task.
+  • Knowledge distillation [82] effectively learns a small model trained to mimic the behavior of a larger, more complex model.
+The knowledge learned by the larger model can be transferred to the smaller model, which can then be used for the target task.
+  • Self-supervised learning [83, 84] involves training a model to predict some aspect of the input data without any external supervision.
+The learned representations can be used for various downstream tasks, including those that involve transferring knowledge from one domain to another.
+3 Time series anomaly detection in industry
+Time series anomaly detection encompasses statistical techniques to analyze and interpret sequential temporal data.
+In the context of industrial processes, time series anomaly detection plays a crucial role in automating monitoring, effectively scheduling maintenance, and controlling the efficiency, quality, and performance of these processes.
+For example, after the detection of an anomaly, another model that captures the relationship between time course and different failure modes or drifts may be exploited for predictive maintenance.
+For example, in injection molding process monitoring, anomaly detection models are used to analyze recorded sensor data from injection molding machines to detect bad parts and identify the root cause of anomalies [85].
+There are two basic ways to detect anomalies: for supervised anomaly detection, labels (normal/abnormal) are needed per time series to build a binary classifier [86].
+For unsupervised anomaly detection, an anomaly score or confidence value that is conditioned purely on normal data can be used to differentiate abnormal from normal instances [87][88].
+3.1 Anomaly types
+According to the literature [89], an outlier is an observation that deviates significantly from other observations in a way that it is likely that it was generated by a different mechanism.
 In this survey, we focus on time series data collected from machine sensor readings in the context of industrial applications, either univariate (only one variable is recorded over time) or multivariate (several simultaneously recorded measurements).
@@
 human error, ambient temperature).
-They can be divided into three categories (Choi et al., 2021; Chalapathy and Chawla, 2019): point anomalies, contextual anomalies, and collective anomalies.
+They can be divided into three categories [32], [33]: point anomalies, contextual anomalies, and collective anomalies.
 Point anomalies are isolated samples that deviate significantly from the normal behavior of that time series, which can be seen on the left of Fig.
-3, e.g., a sudden spike in a pressure reading from a manufacturing machine sensor.
+5, e.g., a sudden spike in a pressure reading from a manufacturing machine sensor.
 These point anomalies can be caused by temporal sensor error, human error, or abnormal machinery operations.
 Contextual anomalies represent data points that deviate from normal ones only in their current context, and an example can be seen in the middle of Fig.
-3.
+5.
 Collective anomalies are a set of data points that in their entirety (but not individually) are abnormal with respect to the entire time series, as shown on the right of Fig.
-3.
+5.
 Figure 5: Three time series anomaly types.
@@
 Point anomalies are single values that fall outside of a pre-set range (left panel).
-Contextual anomalies are samples that deviate from current context (middle panel).
+Contextual anomalies are samples that deviate from the current context (middle panel).
 Collective anomalies are defined as a series of data points that all fall within the range of operation but jointly are not expected (right panel).
-Challenges regarding detecting time series anomalies persist due to two specific properties: First, the complexity of time series data.
+3.2 Challenges
+Challenges regarding detecting time series anomalies persist due to two specific properties: (1) The complexity of time series data.
 As the automation level of industrial processes and the complexity of industrial systems increases, univariate time series data become insufficient and inefficient in representing any industrial process in its entirety.
-Hence, more sensors are installed to monitor the whole process, making it necessary in turn to detect anomalies from multivariate time series, which poses particular challenges since it requires consideration of temporal dependencies and relationships between variables and modalities.
-Second, the dynamic variability of industrial processes can pose difficulties in detecting anomalies due to fluctuations in the process caused by varying input or environmental conditions such as material, temperature, pressure, and humidity, which lead to domain shifts.
-Process automation
-The tasks described above are combined to automate industrial processes.
-For example, after the detection of an anomaly, another model that captures the relationship between time course and different failure modes or drifts may be exploited for predictive maintenance.
-For example, in injection molding process monitoring, anomaly detection models are used to analyze recorded sensor data from injection molding machines to detect bad parts and identify the root cause of anomalies (Tercan et al., 2018).
-There are two basic ways to detect anomalies: for supervised anomaly detection, labels (normal/abnormal) are needed per time series to build a binary classifier (Görnitz et al., 2013).
-For unsupervised anomaly detection, an anomaly score or confidence value that is conditioned purely on normal data can be used to differentiate abnormal from normal instances (Zhang et al., 2023; Audibert et al., 2020; Stadelmann et al., 2019).
+Hence, more sensors are installed to monitor the whole process, making it necessary to detect anomalies from multivariate time series, which poses particular challenges since it requires consideration of temporal dependencies and relationships between variables and modalities.
+Many researchers work on discovering generalized patterns from spatial and temporal correlated multivariate time series data.
+Zhang et al.
+propose a Deep Convolutional Autoencoding Memory network [87], where they build an autoencoder to capture spatial dependency of multi-variant data using MMD to distinguish noisy, normal and abnormal data.
+Zhu et al.
+propose an interpretable model agnostic multivariate time-series anomaly detection method for applications of cyber physical systems [90].
+The new method considers both the temporal and feature dimensions through an adaptive mask based series saliency module to produce accurate anomaly detection results and reasonable interpretations in the form of a mask matrix.
+(2) The dynamic variability in industrial processes.
+Industrial processes often have high dynamic variability and can be affected by a wide range of conditions, such as changes in temperature, pressure, and humidity.
+These conditions can cause fluctuations in the process outputs, which leads to data shift and domain shift.
+This can make it challenging to detect anomalies and maintain control over the industrial process.
+3.3 Anomaly detection methods
+Time series anomaly detection has been investigated for decades, and various types of methods have been proposed [91].
+This paper exclusively discusses the time series anomaly detection techniques using deep learning, leveraging its robust representation learning capabilities.
+Current deep learning methods can be mainly divided into reconstruction-based, forecasting-based, and other methods.
+Fig.
+6 illustrates the two main methods.
+In deep reconstruction-based anomaly detection, the reconstructed sequence is used to compare with the actual sequence.
+Differently, in deep forecasting-based anomaly detection, only the forecasted sequence is used to assess the similarity to the ground truth.
+Figure 6: Illustration of deep learning-based anomaly detection (top row) with reconstruction-based (center row) and forecasting-based (bottom row) anomaly detection in time series.
+The first column represents two time series.
+The second column shows the reconstructed (top) and forecasted (bottom) time series.
+The third column shows the difference between reconstructed/forecasted time series.
+Deviations from the reconstructed or forecasted time series are indicative of an anomaly.
+In deep reconstruction-based anomaly detection, the entire sequence is reconstructed in a decoder-encoder architecture, and the reconstructed sequence is used to compare with the actual sequence.
+In deep forecasting-based anomaly detection, the end of a sequence is predicted using the start of the sequence and only the forecasted sequence is used to assess the similarity to the ground truth.
+In this example, the red time series has a likely anomaly at about 1.5 seconds.
+(Best viewed in color.)
+Reconstruction-based methods
+Reconstruction-based methods aim to learn the data distribution of the normal time series and differentiate the abnormalities from the normal ones by computing the reconstruction errors.
+Audibert et al.
+propose a fast and stable method – Unsupervised anomaly detection for multivariate time series [92], based on adversely trained autoencoders.
+The encoder-decoder architecture within an adversarial training framework combines the advantages of autoencoders and adversarial training while compensating for the limitations of each technique.
+After training two autoencoders, the anomaly score is defined by balancing the reconstructed errors from the two autoencoders with two hyperparameters.
+However, the challenge arises in selecting these two hyperparameters of the anomaly score when the testing dataset is unavailable.
+Malhotra et al.
+also formulate an anomaly score based on reconstruction error [93].
+They first train the LSTM encoder-decoder model to reconstruct the normal time series.
+Subsequently, they leverage the reconstruction errors to calculate the probability by using Maximum Likelihood Estimation to detect a specific point within a time series as an anomaly.
+They set a window to detect anomalies.
+The window will be labeled as anomalous if the probability exceeds a threshold.
+Similarly, Wei et al.
+also propose an LSTM-based encoder-decoder model to detect multivariant time series sequences based on the reconstruction error [94].
+The major difference is the anomaly detection criteria.
+They assume the reconstruction error of train/test data follows the normal distribution and detect anomalies by using the 2-sigma rule of the normal distribution as a threshold.
+Zeng et al.
+propose an adversarial transformer structure to detect multivariate time series anomalies effectively [95].
+Here, two-stage adversarial training is applied for the transformer.
+In the first stage, two transformers are trained by minimizing the reconstruction error to capture the temporal trends in the time series.
+In the second stage, the reconstruction error serves as prior knowledge in the adversarial training process, enabling the model to distinguish anomalies from normal time series.
+Then, an anomaly score is defined by combining the anomaly probability and reconstruction error.
+Again, a threshold has been chosen to differentiate anomalies from normal ones.
+GANs, as effective unsupervised learning methods, have been used in time series anomaly detection.
+Anomaly detection methods based on GANs focus on extracting features by adversarial training on normal samples.
+Consequently, features from the abnormal samples diverge from those of the normal ones, reflecting in reconstruction error and discrimination value.
+Li et al.
+use LSTM-RNN as a base model for building generator and discriminator in GAN[96].
+The proposed framework considers multiple variables to capture the temporal correlation of multi-time series distributions.
+Additionally, they proposed a novel anomaly score, which can detect anomalies through discrimination and reconstruction.
+More specifically, the score is a combination of the reconstruction difference between generated data and original data and the discrimination results from the discriminator.
+Similarly, Niu et al.
+and Bashar et al.
+both propose an LSTM-based VAE-GAN for time series anomaly detection, where LSTM networks are used as the generator, and discriminator [97, 98].
+When it comes to anomaly scores, setting an optimal threshold is usually a critical step.
+However, using a small portion of the test set to decide the optimal threshold may not be practical in real-world scenarios [97].
+Additionally, it is important to note that the method has been only tested for point anomaly detection, further investigation is required when they are applied to detect other anomaly types.
+Forecasting-based methods
+Forecasting-based methods predict the value of the following timestamps and predict temporal anomalies according to the prediction error.
+Kim et al.
+propose a forecasting-based unsupervised time-series anomaly detection method using transformer architecture [99].
+The idea is to train a transformer-like model by forecasting a fixed-length time series based on the previous timestamps.
+The trained model is used to predict time series with an anomaly score such that an instance where the anomaly score is larger than a static threshold is defined as an anomaly.
+A dynamic thresholding technique is also mentioned but not explicitly discussed in the paper.
+Deng and Hooi propose a novel attention-based graph neural network approach [100] that learns a graph of dependence relationships between multi-variant time series signals by forecasting the behavior based on past time series.
+Then, a graph deviation scoring is defined for each sensor to detect and explain anomalies.
+Tang et al.
+propose an interpretable multivariate time series anomaly detection method based on graph neural networks and gated recurrent units [101].
+The feature representation is learned through forecasting the future time series segment.
+An abnormal score is set for each time series to detect anomalies.
+The feature embedding is then used for 2D visualization through t-SNE plots to interpret the clusters within time series from different sensors.
+Other methods
+Ding et al.
+propose a joint network to integrate the advantages of reconstruction and forecasting/prediction [102].
+First, they propose a multimodal graph attention network to tackle the spatial-temporal dependencies for multimodal time series.
+Further, they optimize the reconstruction and prediction modules simultaneously to predict anomalies.
+Himeur et al.
+take advantage of annotated data and directly use a DNN as a classifier to classify normal and abnormal energy consumption types [103].
+The enormous imbalance of real anomaly patterns is one concern in the approach.
+Thus, a normalized technique of power consumption data is applied to deal with this problem.
+The normalized data represent the difference in power consumption rates of each current time sample and the previous one.
+It can provide information on how fast the consumption reacts to the time evolution.
+However, any further evaluation of this technique is not discussed, and it is still an open problem regarding anomaly detection for other datasets.
+Yang et al.
+propose a contrastive learning structure with dual attention to learn a permutation invariant representation of the data with superior discrimination characteristics between normal points and anomalies [104].
+Unlike most reconstruction-based models, their model is a self-supervised framework based on representation learning.
+The new method achieves state-of-the-art comparable performance on six multivariate and one univariate time series anomaly detection benchmark datasets.
+However, the extensive framework with two multi-head-attention blocks may be prone to overfitting.
+This concern is amplified by the absence of training details, leaving only evaluation details disclosed.
+In principle, these anomaly detection approaches are applicable to all types of anomalies.
+Reconstruction-based methods are typically applied to the entire or a portion of the time series.
+Long-time series are commonly segmented into subsequences using a predefined sliding window.
+In the case of detecting context/collective anomalies, the reconstructed loss of the time series sequence is evaluated within the predefined sliding window, if the reconstruction loss is larger than an acceptable threshold, then that time series sequence is classified as an anomaly.
+In the case of point anomalies, reconstruction is performed at each single time stamp, akin to a regression problem, and then the reconstruction loss of each single timestamp is evaluated to determine whether the single timestamp is anomalous or not.
+It is also applied to forecasting-based anomaly detection methods, instead of computing reconstruction error, forecasting-based methods predict the value in the next time stamp for point anomalies or the next time series sequence for context/collective anomalies.
+The anomalies will be detected based on the deviation between the predicted value and the normal value.
+Other anomaly detection approaches usually combine the reconstruction-based and foresting-based methods.
+To sum up, these methods are applicable to each type of anomaly.
+However, the effectiveness of these anomaly detection approaches may vary depending on the anomaly detection tasks at hand, which are characterized by the granularity at which the time series data is observed and analyzed.
 4 Industrial applications
 4.1 Overview
-Currently, deep transfer learning approaches are popular in the field of computer vision and natural language processing because of the large available datasets.
-This popularity is not as pronounced for industrial time series data, likely due to the lack of publicly available data and the domain-specific differences between such data, making the field a less easy target for general gains.
-Fortunately, in recent years a growing number of deep transfer learning approaches have been applied in industry to solve anomaly detection tasks, such as fault diagnosis (Li et al., 2022), quality management (Ma et al., 2019), manufacturing process monitoring (Tercan et al., 2018), network/software security (Rosenberg et al., 2018), and infrastructure monitoring (Pan et al., 2023).
+Deep transfer learning techniques have gained prominence in computer vision and natural language processing, primarily due to the abundance of available datasets.
+However, their adoption in the context of industrial time series data has been comparatively limited.
+This hesitancy can be attributed to the limited public availability of such datasets and the unique domain-specific characteristics they possess, which complicate generalized advancements.
+Encouragingly, there has been a recent uptick in the application of deep transfer learning for anomaly detection within the industry such as fault diagnosis [105], quality management [106], manufacturing process monitoring [85], network/software security [107], and infrastructure monitoring [108].
 These can be mapped onto the core industrial domains of manufacturing process and infrastructure monitoring, predictive maintenance, and energy management.
 Table 2 presents a compact comparison of the related works using deep transfer learning approaches to solve these tasks.
-Figure 4 illustrates the quantity structure of the connections between industrial applications and the deep transfer learning approaches based on our literature survey.
-The Sankey diagram shows every path that connects the four dimensions of the methodology-problem-landscape within the surveyed literature.
-The broader the path is, the more papers are related to that element.
-The goal is to give an overview of how deep transfer learning is applied to industrial problems in the recent literature and specifically show with these four dimensions: (1) which deep transfer learning approaches are actually used in practice; (2) what the main industrial domains for time series anomaly detection are; (3) what deep transfer learning category these domains belong to; and (4) what labels are available in source and target domain.
-Key observations from Figure 4 are: Regarding deep transfer learning approaches, parameter transfer is much more frequently used than any other deep transfer learning approach across all surveyed industrial applications since fine-tuning a pre-trained model on target data is more straightforward to implement by taking advantage of the pre-trained model on the source dataset and usually without fundamental modification on the model architecture.
+Fig.
+7 illustrates the Sankey diagram of the connections between industrial applications and the deep transfer learning approaches based on our literature survey.
+The diagram shows every path that connects the four dimensions of the methodology-problem-landscape within the surveyed literature.
+The broader the path is, the more papers are related to the linked topics.
+The goal is to give an overview of how deep transfer learning is applied to industrial problems in the recent literature and specifically show with these four dimensions: (1) which deep transfer learning approaches are actually used in practice; (2) what the main industrial domains for time series anomaly detection are; (3) what deep transfer learning category these domains belong to; (4) what labels are available in source and target domain.
+Key observations from Fig.
+7 are: (1) parameter transfer is much more frequently used than any other deep transfer learning approach across all surveyed industrial applications since fine-tuning a pre-trained model on target data is more straightforward to implement by taking advantage of the pre-trained model on the source dataset and usually without fundamental modification on the model architecture.
 It is noteworthy that instance transfer and adversarial transfer do not appear in the diagram.
-Apparently, these two deep transfer learning approaches are not considered the optimal choice for respective time series anomaly detection tasks in the industry.
+Apparently, these two deep transfer learning approaches are not considered effective in time series anomaly detection tasks in industry.
 The difficulty lies in implementing and training these scarcely researched approaches in the industrial field, as indicated by the findings.
-Regarding industrial applications, hybrid approaches of parameter and mapping transfer can be seen in predictive maintenance.
-Regarding deep transfer learning categories, most industrial applications use inductive TL, indicating they focus on leveraging labeled source and target data to solve the target task, i.e., use supervised learning.
+(2) Hybrid approaches of parameter and mapping transfer can be seen in predictive maintenance.
+(3) Most industrial applications use inductive transfer learning, indicating they focus on leveraging labeled source and target data to solve the target task, i.e., use supervised learning.
+Figure 7: Overview of Sankey diagram of transfer learning problem setting, deep transfer learning approach categories, and label availability in the surveyed industrial domains.
 Table 2: A compact overview of industrial applications that used deep transfer learning for time series anomaly detection.
@@
 Deep learning framework
-Source data type
+Source type
 Source label
 Target label
-Maschler et al.,2021
+[11]
 Industrial metal forming anomaly detection
@@
 CNN
-Multiple sources
-Wen and Keyes,2019
+Multiple
+[12]
 Monitoring systems Anomaly detection
@@
 U-Net
-Multiple sources
-Xu et al.,2019
+Multiple
+[13]
 Car body-side production line fault diagnosis
@@
 SAE
-Multiple source
-Mao et al.,2020
+Multiple
+[14]
 Rotation bearings fault detection
@@
 Auto-encoder
-Multiple sources
-Wang et al.,2021
+Multiple
+[15]
 Industrial control systems anomaly detection
@@
 ResNet8
-Single source
-Canizo et al.,2019
+Single
+[16]
 Service elevator fault detection
@@
 Inductive
-CNN+RNN
-Multiple sources
-Yao et al.,2022
+CNN, RNN
+Multiple
+[19]
 Nuclear power plants fault detection
@@
 CNN
-Multiple sources
-Li et al.,2023
+Multiple
+[109]
 Building energy systems fault diagnosis
@@
 CNN
-Multiple source
-Serradilla et al.,2021
+Multiple
+[110]
 Press machine production prediction
@@
 CNN
-Single source
-Zgraggen et al.,2021
+Single
+[111]
 Wind turbine fault detection
@@
 CNN
-Single source
-Zabin et al.,2023
+Single
+[112]
 Industrial machine operating fault detection
@@
 Inductive
-CNN+LSTM
-Single source
-Liao et al.,2021
+CNN, LSTM
+Single
+[9]
 Machine turning operations classification
@@
 Inductive
-VGG-19, ResNet
-Single source
-Lockner and Hopmann,2021
+VGG, ResNet
+Single
+[10]
 Injection molding process quality control
@@
 FCN
-Multiple sources
-Tercan et al.,2018
+Multiple
+[85]
 Injection molding process quality control
@@
 FCN
-Single source
-Tercan et al.,2019
+Single
+[113]
 Injection molding process anomaly detection
@@
 FCN
-Multiple sources
-Lockner et al.,2022
+Multiple
+[114]
 Injection molding process anomaly detection
@@
 FCN
-Multiple sources
-Gellrich et al.,2021
+Multiple
+[115]
 Aluminum gravity die casting quality prediction
@@
 FCN
-Single source
-Abdallah et al.,2021
-Digital agriculture and smart manufacturing systems anomaly detection
+Single
+[20]
+Agriculture/manufacturing systems anomaly detection
 Manufacturing process monitoring
@@
 LSTM
-Single source
-Abdallah et al.,2023
+Single
+[116]
 Manufacturing testbeds anomaly detection
@@
 LSTM, RNN
-Single source
-Maschler et al.,2021
+Single
+[117]
 Industrial metal (pump) forming anomaly detection
@@
 LSTM
-Single source
-Panjapornpon et al.,2023
+Single
+[118]
+Industrial control systems anomaly detection
+Manufacturing process monitoring
+Parameter transfer
+Inductive
+Auto-encoder
+Single
+[21]
 Petrochemical production process anomaly detection
@@
 Inductive
-LSTM, CNN, FCN
-Single source
-Liang et al.,2018
-Electricity consumption anomaly detection in aluminium extrusion processes
+LSTM, CNN
+Single
+[119]
+Electricity consumption anomaly detection
 Energy saving
@@
 FCN
-Single source
-Copiaco et al.,2023
+Single
+[120]
 Building’s energy consumption anomaly detection
@@
 Inductive
-MAlexNet-40
-Single source
-Xu et al.,2021
+AlexNet-40
+Single
+[121]
 Power consumption anomaly detection
@@
 Transductive
-Cluster-based DAN
-Single source
-Simone and Amigoni,2021
+DAN
+Single
+[122]
 Building’s power consumption anomaly detection
@@
 LSTM
-Single source
-Xiong et al.,2018
+Single
+[23]
 Aircraft flight anomaly detection
@@
 LSTM
-Single source
-Pan et al.,2023
+Single
+[108]
 Anomaly identification for bridge groups
@@
 CNN
-Single source
-Dhillon and Haque,2020
+Single
+[22]
 Network intrusion detection
@@
 Inductive
-CNN+LSTM
-Single source
-Figure 4: Overview of quantity structure of transfer learning problem setting, deep transfer learning approach categories, and label availability in the surveyed industrial domains.
+CNN, LSTM
+Single
+[17]
+Building occupation detection
+Infrastructure facilities monitoring
+Parameter transfer
+Inductive
+CNN
+Single
+[18]
+Building occupation detection
+Infrastructure facilities monitoring
+Parameter transfer
+Inductive
+CNN
+Single
 4.2 Manufacturing process monitoring
-Manufacturing process monitoring is crucially important to ensure high-quality products and low rejection rates.
-For example, in injection molding machines, sensors are installed to detect molding conditions in the cavity, such as cavity pressure and temperature.
-These signals are used to analyze in particular the mold filling and solidification process for each produced part.
+Manufacturing process monitoring is crucial to ensure high-quality products and low rejection rates.
+For example, in injection molding machines, sensors are installed to detect molding conditions in the cavities, such as cavity pressure and temperature.
+These signals are used to analyze particularly the mold filling and solidification process for each produced part.
 Such cyclic processing data can also be seen in metal machining (cutting force signal) or joining of parts (joining force signal).
-Currently, parameter transfer is predominantly used for manufacturing processes (Lockner and Hopmann, 2021; Tercan et al., 2018, 2019; Lockner et al., 2022; Maschler et al., 2021; Abdallah et al., 2021, 2023; Hsieh et al., 2019; Maschler et al., 2021).
-In injection molding, parameter transfer is applied to transfer the knowledge from one or more source domains to solve tasks in a target domain (Lockner and Hopmann, 2021; Tercan et al., 2019; Lockner et al., 2022).
-Similarly, Tercan et al.
-(2018) build a bridge between simulated data and real data using parameter transfer in injection molding.
+Currently, parameter transfer is predominantly used for manufacturing processes [10, 85, 113, 114, 11, 20, 116, 123, 117].
+Park et al.
+propose a transfer learning technique to detect time series anomalies for different industrial control systems [118].
+First, they apply principal components analysis to reduce the dimension of source and target data.
+A DNN model is then trained on the compressed source data, and after a reasonable mapping algorithm is adopted to map the features of source to target domain, the pre-trained model is further trained on the target data.
+The model achieves good performance even when a model is retrained with only a proportion of target data.
+For the experiments, they only test on two comparatively larger datasets and fail to show that the transferred model performs better than the one without retraining for one dataset.
+Further investigation needs to explain the negative transfer.
+Additionally, even when they only take a small proportion of target data for transfer learning purposes, the sample size still exceeds $5000$, which exceeds most industrial applications.
+Abdallah et al.
+apply parameter transfer to monitor the operation status of manufacturing testbeds with vibration sensor data [20, 116].
+Hsieh et al.
+transfer knowledge across three chambers in a production line to detect anomalous time series data [123].
+Results show reduced training time and improved detection accuracy through transfer learning.
+In injection molding, parameter transfer is applied to transfer the knowledge from one or more source domains to solve tasks in a target domain [10, 113, 114].
+Specifically, they employ simple fully connected neural networks and transfer knowledge from one product to another by freezing the first few layers and fine-tuning only the last few layers.
+Instead of evaluating the time series data directly from sensors, they represent the industrial process by the parameters of the machine settings.
+However, they can still provide useful insight for the case of the time-series data.
+Tercan et al.
+build a bridge between simulated data and real data using parameter transfer in injection molding [85].
+Here, a fully connected neural network is trained on simulated data and then partially or fully reused to further train on real data.
+Results show that the transferred model performs better than a network trained from scratch on real experimental data.
+In manufacturing processes, a simulation model/process hence can play a critical role, but deeper analysis is needed to further understand and reduce the gap between simulation and real data.
+Additionally, Lockner et al.
+explore the impact of transfer learning with varying amounts of source data and assess how performance is influenced by different configurations of frozen layers [114].
 Maschler et al.
-(2021) compare different DNNs for anomaly detection tasks on metal forming datasets.
+compare different DNNs for anomaly detection tasks on metal forming datasets [11].
 Further, they propose a deep transfer learning framework aiming to transfer knowledge between tasks.
-However, the proposed architecture is not validated.
+However, the proposed architecture is not validated by experiments.
 Later, Maschler et al.
-(2021) apply continuous learning on the same dataset by transferring knowledge from several source tasks to a target task to train a deep learning algorithm capable of solving both source and target tasks.
-Abdallah et al.
-(2021, 2023) apply parameter transfer to monitor operation status of manufacturing testbeds with vibration sensor data.
-Hsieh et al.
-(2019) transfer knowledge across three chambers in a production line to detect anomalous time series data.
-Results show reduced training time and improved detection accuracy through transfer learning.
+apply continuous learning on the same dataset by transferring knowledge from several source tasks to a target task to train a deep learning algorithm capable of solving both source and target tasks [117].
+Specifically, they use regularization approaches using altered loss functions to solve related tasks that appear best suited.
 4.3 Predictive maintenance
 Predictive maintenance aims to predict the necessity of maintenance before production is negatively impacted by a failure.
-Tasks involve monitoring equipment to anticipate maintenance requirements (i.e., predict likely future failure) to optimize maintenance schedules (Serradilla et al., 2022).
-Time series anomaly detection is often used in respective systems to identify abnormal patterns or behaviors in operation that may indicate the need for maintenance, such as increasing noise, vibrations, etc.
+Tasks involve monitoring equipment to anticipate maintenance requirements (i.e., predict probable future failure) to optimize maintenance schedules [124].
+Time series anomaly detection is often used in respective systems to identify abnormal behaviors in operation that may indicate the need for maintenance, such as increasing noise, vibrations, etc.
 Mao et al.
-(2020) use mapping transfer with a Sparse Auto-Encoder (SAE) for motor vibration anomaly detection.
-A transformation from the source and target space to a common latent feature space is learned by MMD to make the feature distribution of two domains as identical as possible.
+use mapping transfer with a Sparse Auto-Encoder (SAE) for motor vibration anomaly detection [14].
+A transformation from the source and target data to a common latent feature space is learned with MMD loss to make the feature distribution of two domains as identical as possible.
 Similarly, Wen et al.
-(2019) also used mapping transfer with an SAE architecture for fault detection of rotation bearings, using an MMD regularizer to extract a common feature representation.
-Subsequently, they propose a new MU-Net architecture to deal with multivariate time series anomaly detection tasks (Wen and Keyes, 2019).
-First, they pre-train a U-Net (Ronneberger et al., 2015) on a large time series dataset for an anomaly detection task.
-Then, they propose a new model MU-Net, built upon U-Net, wherein each channel they can use the pre-trained U-Net through fine-tuning to transfer knowledge for multivariate time series anomaly detection.
-In another application, parameter transfer is used to predict the remaining useful life for tools in manufacturing (Sun et al., 2019).
+also use mapping transfer with an SAE architecture for fault detection of rotation bearings, using an MMD regularizer to extract a common feature representation [125].
+Subsequently, they propose a new MU-Net architecture to detect multivariate time series anomalies [12].
+First, they pre-train a U-Net [126] on a large time series dataset for an anomaly detection task.
+Then, they propose a new model MU-Net, which is built upon U-Net.
+In MU-Net, each channel can leverage a pre-trained U-Net through fine-tuning to transfer knowledge for multivariate time series anomaly detection.
+In a different application, parameter transfer is used to predict the remaining useful life for tools in manufacturing [127].
 An SAE network is first trained to predict the remaining useful life of a cutting tool on retrospectively acquired data in an offline process.
-The trained network is then transferred to production with a new tool in operation for online remaining useful life prediction.
-A 2D CNN-LSTM (Hochreiter and Schmidhuber, 1997) hybrid architecture for fault detection is presented.
-The model is trained on a fault dataset, and then parameter transfer is applied to target datasets with a different set of conditions.
-The result shows that transfer learning based hybrid deep learning significantly reduces the training time and is highly suitable for real-time industrial fault diagnosis in various environments.
-Similarly, parameter transfer is implemented to reduce the gap between different industrial environments (Zabin et al., 2023; Xu et al., 2019).
+The trained network is then transferred to production with a new tool for online remaining useful life prediction.
+The result shows that transfer-learning based hybrid deep learning significantly reduces the training time and is highly suitable for real-time industrial fault diagnosis/prediction in various environments.
+Similarly, parameter transfer is implemented to reduce the gap between different industrial environments [112, 13].
 Xu et al.
-(2019) use a stacked SAE to extract general features from source data and a digital-twin-assisted fault diagnosis approach is presented to transfer knowledge from virtual space to physical space for real-time use.
-Here, a DNN model is first fully trained in virtual space and then migrated to the physical space using deep transfer learning for real-time use.
-The surveyed literature proves that deep transfer learning is a research field that could simplify the life cycle of predictive maintenance systems and facilitate DNN model reusability by reducing the required data and training time, helping adapt them to solve similar tasks.
+use a stacked SAE to extract general features from source data and a digital-twin-assisted fault diagnosis approach is presented to transfer knowledge from virtual space to physical space for real-time use [13].
+Here, a DNN model is first fully trained in virtual space and then migrated to the physical space using parameter transfer for real-time use.
+The above-mentioned literature proves that deep transfer learning is a research field that could simplify the life cycle of predictive maintenance systems and facilitate DNN model reusability by reducing the required data and training time, helping adapt them to solve similar tasks.
 4.4 Energy management
-Energy management deals with systems that detect abnormal excessive consumption caused by end-users’ unusual behavior or malfunction of faulty devices or systems (Copiaco et al., 2023).
+Energy management deals with systems that detect abnormal excessive consumption caused by end-users’ unusual behavior or malfunction of faulty devices or systems [120].
 The goal is to develop automatic, quick-responding, accurate, and reliable fault detection to save energy and build environmentally friendly systems.
-Energy anomaly detection systems monitor data during energy generation, transmission, and utilization, in order to ensure normal energy consumption.
+Energy anomaly detection systems monitor data during energy generation, transmission, and utilization, to ensure normal energy consumption.
 Xu et al.
-(2021) design a cluster-based deep adaptation layer to improve a deep adaptation network, effectively reducing the mismatch in transfer learning of spinning power consumption anomaly detection.
+design a cluster-based deep adaptation layer to improve a deep adaptation network, effectively reducing the mismatch in transfer learning of spinning power consumption anomaly detection [121].
+The basic architecture consists of five convolutional layers and three fully connected layers.
+The weight parameters of the convolutional layers are shared between source and target domains.
+The cluster-based deep adaptation layer is designed across the feature layers of two networks to cluster feature representations of the source and target domains respectively.
+The proposed method shows superiority over fine-tuning and DAN because the adaptation layer can minimize the distance between the nearest neighbor clusters across the source and target domains to match the most similar distribution of feature representations.
+It is important to note that the anomalies are defined and tagged by human experts as different types, thus the problem becomes a classification task.
+However, in real-world industrial applications, it’s almost impossible to enumerate unknown anomaly types because of the highly dynamic environment.
 Liang et al.
-(2018) successfully build an electricity consumption time series anomaly detection method in aluminum extrusion.
+successfully build an electricity consumption time series anomaly detection method in aluminum extrusion [119].
 Parameter transfer is applied to transfer domain knowledge from another data-sufficient domain.
-They also find it unnecessary when the target data is sufficient because transferring knowledge decreases prediction accuracy.
+First, they train on sufficient extruding machine data in an unsupervised way and then use only a few data samples from different extruding machines to adapt the model by transfer learning.
+It is important to note that when the target data is already sufficient, transferring knowledge can be detrimental as it can decrease prediction accuracy on the final task.
+Copiaco et al.
+aims to detect anomalies for building energy consumption via transfer learning from pre-trained CNN models [128].
+First, they convert 1D time series signals to 2D image representations.
+These serve as inputs for pre-trained vision models to capture inherent spatially invariant features.
+In the end, a SVM is applied to classify anomaly types.
+The SVM classifier obtained optimal results when operating upon a pre-trained ALexNet model with normalized grayscale graphical representations.
+However, a deeper discussion regarding the effect of the different pre-trained models is not presented.
+Additionally, converting 1D time series to 2D images by creating a matrix representation of the sensor readings may lead to information loss during the transformation process, which should be further investigated.
 4.5 Infrastructure facilities monitoring
-Infrastructure facilities monitoring refers to monitoring and maintaining the conditions of infrastructure facilities, such as bridges, buildings, and networks.
-This can include detecting potential issues or failures.
+Infrastructure facilities monitoring refers to monitoring and maintaining the conditions of infrastructure facilities, such as bridges, buildings[129], and networks.
+This can include detecting potential problems or failures.
 The goal is to minimize the impact of failures on the public or the environment.
+This application commonly uses parameter transfer to transfer knowledge from facility to facility to take advantage of similar data and tasks.
+Dhillon et al.
+present a parameter transfer approach towards building a network intrusion detection system based on CNN and LSTM [22].
+Specifically, They extract and learn patterns by mapping the input data into a lower dimensional representation by convolutional layers.
+Then, they employ the LSTM layer to enhance learning and recognizing patterns across time.
+In the end, a fully connected layer is used as a classifier to predict normal and malicious data.
+To do the parameter transfer, they reuse the model architecture and freeze most weight parameters for the target domain so that they do not need a large training dataset to retrain the model.
+However, they do not mention implementation details, like which layers are frozen in the transfer learning stage.
+Observing how transfer learning performs with different frozen layers would be interesting.
 Pan et al.
-(2023) apply parameter transfer to make full use of the similarity of the anomalous patterns across different bridges and transfer the knowledge obtained by a CNN model to a small part of target data, achieving high accuracy anomaly detection across bridges.
-Dhillon and Haque (2020) present a parameter transfer approach towards building a network intrusion detection system based on CNN and LSTM.
+apply parameter transfer to fully use the similarity of the anomalous patterns across different bridges [108].
+They train a CNN model on one bridge data, then transfer the knowledge obtained by the CNN model to a small part of the target data.
+They update the last three fully connected layers while keeping the convolutional layers intact.
+The experimental results show transfer learning achieves higher accuracy anomaly detection across bridges.
+Weber et al.
+takes advantage of simulation data by training on synthetic environmental data, then fine-tunes the pre-trained model and transfers the knowledge from simulation data for real-time online building occupancy detection [17].
+Although the results show the effectiveness of transfer learning, the availability, and reliability of the simulation data for other industrial applications is still an open issue.
+Sayed et al.
+adapt parameter transfer using pre-trained CNN models, such as AlexNet and GoogLeNet, pre-trained on ImageNet [18].
+The pre-trained model is then further used for downstream tasks.
+The results show the pre-trained models outperform their customed CNN model, which is not pre-trained.
+However, it is important to note that the transfer effect may not be entirely convincing due to the fact that the customer CNN model is not pre-trained on the same dataset.
+4.6 Application-independent considerations
+[…]
+Data scarcity, as well as domain shift, stand out as the two main common problems independent of the industrial field of application.
+The same problems have originally prompted the use of transfer learning in general, and respectively, general techniques are applied widely across domains.
+Regarding data scarcity, this un-surprisingly involves leveraging pre-trained models as a starting point for further training.
+Regarding domain shift, mapping transfer and parameter transfer are the most often-used approaches.
+Unlike parameter transfer, mapping transfer incorporates the source and target data in the training process.
+Instance transfer and domain-adversarial transfer learning were not employed in the surveyed literature – researchers seem to not see huge value in these methods for the surveyed fields.
+Another common aspect across time series anomaly detection applications is the choices of model architecture to facilitate the training process by capturing temporal dependencies and recognizing patterns over varying time scales: Favourite architectures include CNNs, LSTMs, and auto-encoders that have sets of assumptions (inductive biases) about the data they analyze that make them excel in understanding the sequential nature of time series data.
+CNNs assume local (in time) connectivity, stationary statistics, and hierarchical structure, and induce certain translation-invariance.
+LSTMs are still given preference in many applications over the more modern deep learning architecture of choice for sequence learning, the transformer.
+The reason is their stronger inductive bias, leading to less data (and compute) hunger.
+Both CNN and LSTM networks can be built as e.g.
+classifiers, but also auto-encoders.
+These latter architectures have the advantage of learning low-dimensional representations of the high-dimensional time series with minimal loss of signal in an unsupervised way.
+The analysis of the data in the low-dimensional latent space facilitates anomaly detection.
+The concrete choice of architecture does not depend on the field of application but on the data and task, and thus the most suitable inductive bias.
+As an interim conclusion, the most striking application-independent finding is that across the surveyed literature, predominantly simple, tried-and-tested design patterns for transfer learning are used in industry.
+The field of deep transfer learning would offer a much wider variety of approaches.
 5 Discussion
+5.1 Potential
+The automation of industrial process monitoring stands as a transformative step toward increasing efficiency and optimizing quality.
+While standard deep learning training is sufficient in discerning intricate patterns from vast datasets, its application in the dynamic industrial landscape is not without challenges.
+Chief among them is the impracticality of continuously obtaining large-scale labeled data to train models afresh for every nuanced variation in processes.
+Deep transfer learning has shown promise with its adaptive capabilities.
+By mitigating the need for extensive labeled data and eliminating the necessity to train models from scratch for every distinct setup.
+However, adopting deep transfer learning beyond simple parameter transfer is still a challenge.
 5.2 Challenges
-Label availability
-Deep transfer learning is built upon deep learning, which usually requires a large amount of labeled data, the more data a model has available for learning, the better it can generalize to new examples.
+Domain shift
+Different from the i.i.d assumption in most machine learning problem settings, many industrial processes suffer from substantial domain shift due to dynamic changes in industrial settings, e.g., change of products or measuring sensors.
+Domain shift lies at the heart of the deep transfer learning problem.
+Particularly, the dynamic changes in many industrial processes, up to an apparent dissimilarity of source and target data, make the transfer learning task particularly challenging.
+Here we list the key challenges associated with domain shift:
+  • Covariate shift occurs when the marginal distribution of the features changes from source domain to target domain.
+The distribution mismatch poses challenges in transferring the knowledge from source to target domain.
+  • Concept shift refers to the changes in the relationships between features and labels. The relationship can change from source domain to target domain, leading to bias and error in the model.
+  • Label shift refers to the label distribution in the target domain that can be different from the source domain, whether the marginal distribution changes or not.
+Label availability and reliability
+Deep transfer learning is built upon deep learning, which usually requires a large amount of labeled data, the more data a model has available for training, the better it can generalize to new examples.
 In real-world industrial time series anomaly detection tasks, collecting data is probably easy, but collecting labels is much more expensive and time-consuming, sometimes prohibitively so, leading to the unavailability of sufficient labeled data.
-Self-supervised learning can be used to re-label a large amount of unlabeled data and thus facilitate transfer learning process.
-Thus, anomaly detection models usually need to learn in an unsupervised or semi-supervised mode (Goldstein and Uchida, 2016).
-Deep learning for imbalanced data
-Even if the labels can be collected, anomalies can be extremely rare by design, which poses the risk of training with extremely imbalanced data.
+Self-supervised learning can be used to re-label a large amount of unlabeled data and thus anomaly detection models usually need to learn in an unsupervised or semi-supervised mode [130].
+In industrial cases, another significant concern is to ensure the data quality.
+Due to the high cost associated with obtaining reliable and precise labels, usually self-supervised learning is applied to create pseudo labels or relabel the unlabeled data, thus facilitating the transfer learning process.
+Additionally, a data-centric process with humans in the loop can be involved in improving label reliability.
+However, unreliable labels can still affect the transfer learning training process.
+Missing relevant data information
+Missing relevant data poses significant challenges for transfer learning since it can affect the model’s ability to generalize and transfer knowledge from source to target domain.
+  • Imbalanced data: Even if the labels can be collected, anomalies can be extremely rare by design, which poses the risk of training with extremely imbalanced data.
 A practical problem for anomaly detection in industry is the extremely imbalanced data distribution, in which normal samples dominate in data and abnormal samples only share a small percentage in the whole dataset.
-Prior research has proven that the effect of class imbalance on classification performance by using deep learning is detrimental (Buda et al., 2018).
+Prior research has proven that the effect of class imbalance on classification performance by using deep learning is detrimental [131].
 However, most research studies still ignore such problems, which can result in poor performance regarding the minority class, i.e., abnormal data are misclassified as normal.
-Missing relevant data
-Another problem is missing relevant data, i.e, some information that has a significant effect on the process from case to case is not even recorded or is too complex to record (i.e.
+  • Information loss: missing data can lead to lost important features.
+For example, some information that has a significant effect on the process from case to case is not even recorded or is too complex to record (i.e.
 part geometry, machine geometry, or environmental conditions in injection molding processes).
-Domain shift
-Domain shifts lie at the heart of the deep transfer learning problem, but the dynamic changes in many industrial processes, up to an apparent dissimilarity of source and target data, make the transfer learning task particularly challenging.
+Various approaches have been developed to address these challenges to reduce the domain gap between the source and target domains, aiming at mitigating domain shift.
+These techniques involve domain generalization, contrastive learning, and adversarial examples.
+The domain shift problem is far from being solved.
+To tackle this problem, transfer learning requires a deep understanding of the target data’s characteristics and appropriate transferable strategies to effectively bridge the gap between the source and target domains.
 Effectiveness of deep transfer learning
 The general effectiveness of deep transfer learning is limited by the difficulty of determining which knowledge or to what extent the knowledge should be transferred from source to target task.
-Unlike natural language processing, pretraining a language model on a large corpus of text data can help the model learn the statistical patterns and semantic and syntactic representations of words and sentences, which can be used for new natural language processing tasks with a few data.
-For industrial time series, due to data privacy, large available public datasets usually do not exist, or they cannot be used even because of a large domain gap between different datasets and tasks.
+Unlike natural language processing, pre-training a language model on a large corpus of text data can help the model learn the statistical patterns and semantic and syntactic representations of words and sentences, which can be used for new natural language processing tasks with a few or even without data.
+Due to data privacy, large available public datasets usually do not exist for industrial time series, or they can not be used because of a large domain gap between different datasets and tasks.
 In this case, transferring all of the knowledge may not be beneficial, as it may be irrelevant.
-In the worst case, this can lead to negative transfer (Pan and Yang, 2010; Smith-Jentsch et al., 2001), in which the extracted knowledge harms the new task-learning.
+In the worst case, this can lead to negative transfer [28, 132], in which the extracted knowledge harms the new task-learning.
 This requires assessing how source and target tasks are related, carefully selecting the knowledge to be transferred, and selecting the proper means to implement this transfer.
+Glorot et al.
+attempt to analyze and quantify the gained knowledge from source to target domain [133].
+For example, they define transfer error, transfer loss, transfer ratio, and in-domain ratio, which provide metrics to interpret the transferring performance.
 5.3 Directions for anomaly detection solution design
@@
 How data preprocessing should be conducted is an open question.
-For industrial applications, some researchers consider directly using time series data as input for training to be inefficient, thus suggesting deriving or selecting features from time series data by statistical methods or human experience to decrease the complexity of the dataset dramatically.
+For industrial applications, some researchers contend that using raw time series data directly as input for training may not be the most efficient.
+Hence, they propose deriving or selecting features from time series data by statistical methods or human experience.
+This can significantly decrease the complexity of the dataset.
 On the other hand, this crops a lot of potentially useful information, e.g., the time series trend.
-Some researchers use machine parameters as features of the manufacturing process instead of using process data collected by sensors (Lockner and Hopmann, 2021; Tercan et al., 2018, 2019; Lockner et al., 2022).
-Others try different transformations of raw time series data, a common way being to transform 1D time series data to 2D image data (Liao et al., 2021; Wang et al., 2021; Zabin et al., 2023) or transforming time domain signals otherwise into the frequency domain (Lukic et al., 2016).
-However, as large-scale computation power and storage become cheaper and more accessible, it is becoming increasingly common to use deep learning techniques to directly process time series data (Xu et al., 2019; Maschler et al., 2021).
-Data augmentation
+To reduce the dimensionality, some researchers use machine parameters as features in the manufacturing process instead of the processing data collected by sensors [85, 10, 113, 114].
+Others try different transformations of raw time series data, a common way being to transform 1D time series data to 2D image data [9, 15, 112] or transforming time domain signals otherwise into the frequency domain [39].
+However, as large-scale computation power and storage become cheaper and more accessible, it is becoming increasingly common to use deep learning techniques to process time series data directly [13, 11].
+Data augmentation by generative AI
 Data augmentation is useful for deep learning models because it can help to prevent overfitting.
 For deep transfer learning, when a model becomes too closely adapted to the specifics of the source domain, it may not be able to generalize well to some examples in the task domain.
-One important technique is to acquire effective synthetic data, e.g., using a simulation process or model to explore potential anomalous conditions by simulating industrial processes under parameters that cannot yet be experienced in the real-world.
-High fidelity and reliability of simulation data can provide training data at low cost and mitigate the problem of insufficient samples for deep transfer learning (Xu et al., 2019).
-Another way to generate effective synthetic data is to use GANs.
-GANs are trained on normal data only to generate indistinguishable samples from which abnormal samples are distinguished during the testing stage of the overall anomaly detection system based on their deviating data distribution (Jiang et al., 2019).
-To increase the number of anomalous samples and thus the robustness of the anomaly detection model, the technique of adversarial perturbation known from computer vision (Goodfellow et al., 2015) can be used.
-Data imbalance
+One important technique is to acquire effective synthetic data, e.g., using a simulation process or model to explore potential anomalous conditions by simulating industrial processes under parameters that cannot yet be experienced in the real world.
+High fidelity and reliable simulation data can provide training data at low cost and mitigate the problem of insufficient samples for deep transfer learning [13].
+Another way to generate effective synthetic data is to use generative models, such as GANs.
+GANs are only trained on normal data to generate indistinguishable normal samples so that abnormal samples can be distinguished during the testing stage of the overarching anomaly detection system, as they deviate from the normal data distribution [134].
+To increase the number of anomalous samples and thus the robustness of the anomaly detection model, the technique of adversarial perturbation known from computer vision [135] can be used.
+Dealing with data imbalance
 DNNs perform well when they are trained on balanced datasets.
 However, in practice, it is difficult to get sufficient anomalous data for anomaly detection tasks.
-For example, manufacturing process is usually in a healthy state due to the pre-designed and optimized operation.
+For example, the manufacturing process is usually in a healthy state due to the pre-designed and optimized operation.
 Several ways exist to address the imbalanced dataset for time series anomaly detection.
-To deal with data imbalance, one way is to oversample the minority class, i.e., to randomly replicate samples from the minority class to equalize the number of samples from each class in each batch.
-Synthetic Minority Over-sampling Technique is an advanced method that creates synthetic samples to force the decision region of the minority class to become more general (Chawla et al., 2002).
-This technique is widely used in anomaly detection tasks in industry (Ijaz et al., 2018; Mokhtari et al., 2021).
-Apart from oversampling, the resampling strategy is also frequently used to assign a higher probability to abnormal samples and evenly select the same amount of samples from both classes in each batch.
-Finally, a weighted loss can be used that balances the loss for the abnormal and normal class in supervised anomaly detection (Buda et al., 2018).
+One way is to oversample the minority class, e.g., by randomly replicating samples from the minority class to equalize the number of samples from each class in each batch.
+The Synthetic Minority Over-sampling Technique is an advanced method that creates synthetic samples to force the decision region of the minority class to become more general [136].
+This technique is widely used in anomaly detection tasks in industry [137, 138].
+Apart from oversampling, resampling strategies are frequently used to assign a higher probability to abnormal samples and evenly select the same amount of samples from both classes in each batch.
+Moreover, a weighted loss can be implemented to balance the loss between the abnormal and normal class in supervised anomaly detection [131].
 5.4 Directions for deep transfer learning implementation
 When shall deep transfer learning be used?
-(1) Limited data availability: If the amount of data available for a specific task is limited, pre-training on related source data can learn general features that can be transferred to the specific learning task in the target domain.
-(2) Similar domains: deep transfer learning is well suited for similar source and target domains.
-(3) Limited resources (time and compute): Using parameter transfer here is recommended if pre-trained models exist.
+[…]
 When not to use deep transfer learning?
-(1) Different tasks: If the target learning task is vastly different from the source learning task, deep transfer learning may not be appropriate.
-For example, if one wants to train a model for natural language processing on a new dataset, using a pre-trained model that has been trained on image data will not be useful.
-(2) High domain shift: If there is a large difference between the source and the target domain, deep transfer learning may not be effective.
-This can happen when the data distributions, features, or labels are vastly different.
-(3) Abundance of labeled data: If there are enough data for the new task, it may be more effective to train a model from scratch (Liang et al., 2018).
+[…]
 What model architecture to choose?
-We suggest selecting the model architecture mainly depending on the data size and label availability, starting from a relatively small network and moving gradually to more complex DNNs.
-It is important to effectively capture the temporal dependencies and extract respective features of time series data.
-LSTMs are used heavily for detecting temporal dependencies in time series data (Zabin et al., 2023; Panjapornpon et al., 2023; Xiong et al., 2018).
-CNNs are also effective in extracting time series features (Yao et al., 2022; Zgraggen et al., 2021).
-For semi-supervised settings, CNN-based auto-encoders are trained to reconstruct the original data (Serradilla et al., 2021).
-Another aspect of deep transfer learning implementations is often the limited computing power of hardware platforms, such as embedded systems in industrial applications.
-Sensor data are typically acquired using resource-constrained edge processing devices that struggle with computationally intensive tasks, especially when updating a DNN model.
-One possible solution is federated learning, perhaps the most popular framework, mainly due to its feature of leveraging data while still preserving their privacy (Wang et al., 2018; Amiri and Gündüz, 2019).
+We recommend choosing the model architecture mainly based on data size and label availability, starting from a relatively small network and moving gradually to more complex DNNs.
+CNNs also effectively extract time series features [19, 111].
+For semi-supervised settings, CNN-based auto-encoders are trained to reconstruct the original data [110].
+It is important to effectively capture the temporal dependencies and extract features of time series data.
+LSTMs are extensively employed for this purpose, as they excel in detecting temporal dependencies in time series data [112, 21, 23].
+[…]
+[…]
+Beyond transfer learning
+Foundation models like SAM [139] or others, using for example transformer architectures [48] or diffusion models [140], demonstrate emerging properties such as in-context learning [53] and complex cross-modality conditioning.
+This is achieved by training complex and often auto-regressive models with massive amounts of data, although the precise mechanisms that lead to this are not well understood.
+Some of those models generalize to new settings and tasks, without an explicit element of transfer learning.
+Thus, the application of foundation models in industrial time series analysis has the potential to reduce and eventually eliminate the need to explicitly account for changes in the domain within the modeling, by instead having the foundation model provide the transfer capability (see examples in Sec.
+6-b).
+To not only detect anomalies but also identify failure modes, analyze root causes, and elicit an appropriate intervention, AI systems must implicitly or explicitly model causal relations.
+Counterfactual inference incorporates causal relations between observations and interventions, which allows predictions of outcomes never seen during training [141].
+Another aspect of deep learning implementations is the limited computing power of hardware platforms, such as embedded systems in industry.
+Sensor data are typically acquired using resource-constrained edge processing devices that struggle with computationally intensive tasks, especially when training a DNN model.
+Federated learning stands out as a leading solution, with its ability to utilize data while preserving privacy [142, 143].
 The technology enables a more collaborative approach to ML while preserving user privacy by storing data decentralized on distributed devices rather than on a central server.
 Combining deep transfer learning with federated learning is a promising and powerful combination in the abovementioned industrial applications.
-Beyond transfer learning
-Foundation models such as transformers (Vaswani et al., 2017), diffusion models (Rombach et al., 2022) and SAM (Kirillov et al., 2023) demonstrate emerging properties such as in-context learning (Brown et al., 2020) and complex cross-modality conditioning.
-This is achieved by training complex and often auto-regressive models with massive amounts of data, although the precise mechanisms that lead to this are not well understood.
-Some of those models generalize to new settings and tasks, without an explicit element of transfer learning.
-Thus, the application of foundational models in industrial time series analysis has the potential to reduce and eventually eliminate the need to explicitly account for changes in the domain in modeling, and instead, the foundational models will provide the transfer capability.
-To not only detect anomalies but also identify failure modes and elicit an appropriate intervention, AI systems must have some form of understanding, a world model, or, in other words, the AI has to implicitly or explicitly model causal relations.
-Counterfactual inference incorporates causal relations between observations and interventions which allows making predictions of outcomes that were never seen during training (Vlontzos et al., 2023).
 6 Conclusions
-In this survey, we presented a comprehensive overview of (deep) transfer learning by defining transfer learning problem settings and categorizing the state-of-the-art deep transfer learning approaches.
+In this survey, we presented a comprehensive overview of deep transfer learning by defining transfer learning problem settings and categorizing the state-of-the-art deep transfer learning approaches based on the surveyed papers.
+Then, we review and emphasize on investigating deep transfer learning approaches for time series anomaly detection in different industrial settings.
 Equipped with this foundation, we selected representative examples of the landscape of fielded applications to provide practitioners with a guide to the field and possibilities of industrial time series anomaly detection.
+The main finding of this survey is that only a limited variety of deep transfer learning methods are employed in anomaly detection in industrial time series analysis – mainly simple ones.
+Almost all applications employ parameter transfer, arguably the most straightforward transfer approach.
+In its simplest implementation, it only involves fine-tuning a pre-trained model.
+Accordingly, the employed network architectures are simple, none of the reviewed research papers used advanced DNN building blocks like Transformer, which are common in computer vision and language modeling.
+We expect this type of architecture with suitable modifications and/or pre-trained parameters to spread to more niche fields.
+Despite this, the survey suggests that deep transfer learning approaches have huge potential and promise for solving more complex and dynamic anomaly detection tasks in industry.
+As the field is still in an early stage, more R&D is expected to fully realize the potential of deep transfer learning in increasingly complex settings.
+In the end, we highlight the importance of considering feasibility, reliability, explainability, and real-time data stream when designing a transfer learning system for time series anomaly detection.
 After carefully discussing open challenges, we gave practical directions for time series anomaly detection solution design and deep transfer learning implementation.
-We found that current applications focus on simple cases with simple datasets, neural network structures, and deep transfer learning schemes.
-Despite this, the survey suggests that deep transfer learning approaches have huge potential and promise for solving more complex and dynamic anomaly detection tasks in the industry.
-As the field is still in an early stage, more R&D is expected to fully realize the potential of deep transfer learning in increasingly complex settings.
-Future work should focus on developing robust transfer learning schemas and methods that can handle more complex and dynamic tasks.
-The following directions hold the greatest potential for future work:
-Automatic selection of transferable features
-(Long et al., 2015) It refers to methods for selecting and transferring only the relevant knowledge from the base model.
-This could involve the use of techniques such as selective fine-tuning and distillation to identify the most important features and knowledge learned from source domains (Yosinski et al., 2014; Ge and Yu, 2017).
-Investing into more complex deep transfer learning schemes and DNN architectures
-Most deep transfer learning approaches applied in industry focus on the parameter transfer approach as it is conceptionally the simplest and readily applicable by interdisciplinary teams without ML research experience.
-It seems promising to invest in testing more appropriate deep transfer learning approaches according to different use cases, such as mapping transfer, adversarial transfer, etc.
-The same applies to testing diverse DNN architectures besides straightforward ones.
+In our view, the following directions hold the greatest potential for future work:
+Automatic selection of transferable features [57]
+It refers to methods for selecting and transferring only the relevant knowledge for the new tasks from the base model.
+This could involve the use of techniques such as selective fine-tuning and distillation to identify the most important features learned from source domains [30, 144].
+Investing into advanced deep transfer learning schemes and DNN models
+[…]
 Data-centric approach to real-time anomaly detection
-The data-centric approach focuses on improving ML models by ensuring high-quality labeled data (Stadelmann et al., 2022) using techniques such as re-labeling, re-weighting, or data augmentation (Luley et al., 2023).
-Currently, a human-in-the-loop solution is still needed, frameworks have been proposed to assist annotators with graph-based algorithms such as nearest neighbor graphs (Bai et al., 2021), decision trees (Liu et al., 2021), or factor graphs (Kang et al., 2022).
+The data-centric approach focuses on improving ML models by ensuring high-quality labeled data [147] using techniques such as re-labeling, re-weighting, or data augmentation [148].
+Currently, a human-in-the-loop solution is still needed.
+Frameworks have been proposed to assist annotators with graph-based algorithms such as nearest neighbor graphs [84], decision trees [149], or factor graphs [150].
 Although these methods have proven to be effective, a more automated process is a goal for future research.
+Leveraging generative AI
+Generative models like GANs and diffusion models can generate synthetic time series data, making them valuable for data augmentation.
+Augmenting the original data with synthetic samples can enhance the deep learning models’ robustness, especially in real-world applications where target data are limited.
+These models can also be leveraged to examine anomalies and generate anomalies to help alleviate the imbalance within the data [151].
 Integration with other ML methods
-To build robust AI approaches to solve time series anomaly detection in industry, only focusing on transfer learning will not be sufficient.
-Combinations with other ML approaches are needed in the future, such as continuous learning, meta-learning, and federated learning.
-Acknowledgments
-We would like to acknowledge Claudio Riginio for his constructive comments on an earlier draft of this paper and for his assistance with the illustrations.
-This work has been supported by Innosuisse grant 62174.1 IP-ENG “DISTRAL”.
+To develop robust AI solutions for time series anomaly detection in the industry, relying solely on transfer learning is insufficient.
+Future strategies should integrate other ML approaches, including continuous learning, meta-learning, and federated learning.